### v3

# Convolution estimates for the Boltzmann gain operator with hard-spheres

Abstract.
We prove new moment-preserving polynomially weighted convolution estimates for the gain operator of the Boltzmann equation with hard potentials, including the critical case of hard-spheres. Our approach relies crucially on a novel cancellation mechanism dealing with the pathological case of energy-absorbing collisions (that is, collisions that accumulate energy to only one of the outgoing particles). This difficulty is specific to hard potentials, and is not present for Maxwell molecules. Our method quantifies the heuristic that, while energy-absorbing collisions occur with non-trivial probability, they are statistically rare, and therefore do not affect the overall averaging behavior of the gain operator. At the technical level, our proof relies solely on tools from kinetic theory, such as geometric identities and angular averaging.
1. Introduction
1.1. Problem setup and goal of this paper
The Boltzmann equation is the central equation of collisional kinetic theory and one of the most celebrated equations of mathematical physics. It describes dilute gases of microscopic interacting particles at the mesoscopic/kinetic level; more precisely, it predicts the time evolution of the probability density $f(t,x,v)$ of a dilute gas in non-equilibrium. The Boltzmann equation reads
(1.1) $\partial_{t}f+v\cdot\nabla_{x}f=Q(f,f),$
where the advection operator $\partial_{t}+v\cdot\nabla_{x}$ accounts for the rectilinear motion of particles, while $Q(f,f)$ is the quadratic collisional operator encoding the binary interactions between gas particles. The collisional operator $Q(f,f)$ is given by
(1.2) $Q(f,f)(t,x,v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}B(u,\sigma)\Big{(}f(t,x,v^{*})f(t,x,v_{1}^{*})-f(t,x,v)f(t,x,v_{1})\Big{)}\,d\sigma\,dv_{1},$
where $v,v_{1}$ denote the pre-collisional velocities of the incoming particles, $v^{*}$, $v_{1}^{*}$ denote the post-collisional velocities, and $u:=v-v_{1}$ denotes the relative velocity of the particles before the collision. Simplifying the notation slightly, we drop the dependence of the arguments on $t,x$ since it is clearly implied by the context.
In the case of perfectly elastic collisions, and assuming without loss of generality that all particles are of equal mass $m=1$, the post-collisional velocities are given by:
(1.3) $\begin{split}v^{*}&=V-\frac{|u|}{2}\sigma,\quad v_{1}^{*}=V+\frac{|u|}{2}\sigma,\end{split}$
where $V:=\frac{v+v_{1}}{2}$ is the center of mass of the velocities of the incoming velocities and $\sigma\in{\mathbb{S}}^{2}$ represents the scattering direction of the collision. It is then straightforward to see that the momentum and energy are conserved, as well as the relative velocity magnitude:
(1.4) $v^{*}+v_{1}^{*}$ $=v+v_{1},$
(1.5) $|v^{*}|^{2}+|v_{1}^{*}|^{2}$ $=|v|^{2}+|v_{1}|^{2},$
(1.6) $|v^{*}-v_{1}^{*}|$ $=|v-v_{1}|.$
The kernel $B(u,\sigma)$ appearing in (1.2) is the collisional cross-section, which is assumed to be of the form
(1.7) $B(u,\sigma)=|u|^{\gamma}\,b(\widehat{u}\cdot\sigma),\quad-3<\gamma\leq 1,$
where the angular cross-section $b:[-1,1]\to{\mathbb{R}}$ satisfies:
(H.1) $b(z)$ $\geq 0,\quad\forall\,z\in[-1,1],$
(H.2) $b(-z)$ $=b(z),\quad\forall\,z\in[-1,1],$
(H.3) $b\in L^{\infty}([-1,1]).$
The parameter $\gamma$ in the radial part of the cross-section is of crucial importance: it determines the type of intramolecular potential considered.
  • The case $0<\gamma\leq 1$ corresponds to hard potentials, where the interaction is strong and of short range. They include the celebrated hard-sphere model ($\gamma=1,\,b=1$), where particles interact like impenetrable billiard balls. Hard-spheres are of particular importance since it was the model originally introduced by Boltzmann [8] and Maxwell [30] and remains the only model for which the Boltzmann has been rigorously derived from deterministic many particle dynamics [27, 22, 15].
  • The case $\gamma=0$ corresponds to Maxwell molecules.
  • The case $3<\gamma<0$ corresponds to soft potentials. This type of interaction is smoother and of longer range.
Both physically and mathematically, these three types of potentials lead to different kind of obstacles that typically require separate treatment. In this paper, we focus on hard potentials. In the course of the proof, we also improve existing results on Maxwell molecules.
Regarding the assumptions on the angular cross-section $b$, condition (H.2) corresponds to indistiguishability of the particles, while the boundedness condition (H.3) is a slightly stronger assumption than the usual Grad’s cut-off, which is the mere integrability condition $\int_{{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)\,d\sigma<\infty.$ Yet (H.3) covers most cases of physical interest for hard potentials, and in particular the classical hard-sphere model ($\gamma=1,\,b=1$). Most importantly, since (H.3) is a special case of Grad’s cut-off, it implies that the collisional operator can be written in gain-loss form
(1.8) $Q(f,f)=Q^{+}(f,f)-Q^{-}(f,f),$
where the gain and loss operators are respectively given by
(1.9) $Q^{+}(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\,d\sigma\,dv_{1},$
(1.10) $Q^{-}(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)f(v_{1})\,d\sigma\,dv_{1}.$
Half-sphere restriction and bilinear operators
Exploiting spherical symmetry, it is often convenient to restrict the collisional operators to the half-sphere. Namely, denoting $\chi:=\mathds{1}_{(0,+\infty)},$ one obtains
$Q(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(-\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(-\widehat{u}\cdot\sigma)\Big{(}f(v_{1}^{*})f(v^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
where in the negative part we used the substitution $\sigma\to-\sigma$ (which just interchanges $v^{*}$ and $v_{1}^{*}$ by (1.3)) and for the last equality we used (H.2).
Motivated by this observation, we extend the collisional operators to bilinear forms in the following way:
(1.11) $Q(f,g)$ $:=Q^{+}(f,g)-Q^{-}(f,g),$
(1.12) $Q^{+}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
(1.13) $Q^{-}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)g(v_{1})\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Clearly, the loss operator $Q^{-}(f,g)$ is much simpler to analyze, and essentially behaves like the product $f(|v|^{\gamma}\ast g).$
In contrast, the gain operator $Q^{+}(f,g)$ has a much richer structure due to the averaging over the angular variable, and exhibits regularization effects [25, 28]. As a result, it is the main object to understand when developing the mathematical theory of the Boltzmann and related equations.
Quantifying these regularizing effects is a subtle task though due to the complex structure of the operator. The first step is to choose a suitable topology to measure the gain of regularity. The most relevant functional analytic setting is polynomially weighted Lebesgue spaces only in the velocity variable ${\langle}v{\rangle}^{-k}L^{p}_{v}$ for several reasons:
  • Such bounds are space invariant thus they can be applied whether the spatial domain is bounded or unbounded. Indeed, ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$ estimates easily follow by Hölder’s inequality. This feature is important to address questions related to the hydrodynamic limit of (1.1). In that field, most mathematical results are proved in spatially bounded domains such as the torus. Without being exhaustive, we refer to the recent work of Deng-Hani-Ma [16] for a complete derivation of fundamental fluid dynamics equations from many-body classical dynamics, which followed their groundbreaking work [15] on the long-time derivation of the Boltzmann equation from hard-sphere dynamics, that extended Lanford’s theorem [27, 22] globally in time.
  • Polynomial weights are the most suitable for studying the long-time behavior of the equation for large data, which is a major open problem, as outlined in [37]. Indeed, large data results for the space homogeneous (i.e. space independent) Boltzmann equation for hard potentials have been proved by establishing generation and propagation of moments [34, 20, 17, 38, 40, 11, 31], which correspond exactly to Lebesgue spaces with polynomial weights in $v$. This position was already advocated by L. Desvillettes and C. Villani [18], where the asymptotic behavior of solutions that remain bounded in such spaces is fully characterized. As noted by these authors, this setting is also outside of the scope of linearization techniques, which typically require exponential weights.
  • Working with polynomial weights is also more natural in the closely related field of wave turbulence, where equilibria are typically inverse polynomials in $v$ (e.g. Rayleigh-Jeans distributions or the celebrated Kolmogorov-Zakharov spectra, see e.g. Nazarenko [33] for more details). Therefore, ideas from this setting could be helpful in studying wave kinetic equations as well, as shown in our previous work [3]. In an upcoming work [5] we are planning to use ideas and techniques of the current paper in order to prove optimal local well-posedness of the wave kinetic equation in $L^{p}$ spaces for any $2\leq p\leq\infty$. Such a result would extend the work of Germain-Ionescu-Tran [23] who obtained optimal well-posedness in $L^{2}$ and in $L^{\infty}$.
  • Finally, the conditions for the solution to be derivable from hard-sphere dynamics, and thus solve Hilbert’s sixth problem require weights only in $v$, see the statement of Theorem 1 in the seminal work of Deng-Hani-Ma [15].
In this paper, we prove new moment-preserving polynomially weighted convolution estimates for $Q^{+}(f,g)$ with hard potentials, including hard-spheres. Our approach relies crucially on a novel cancellation mechanism dealing with pathological collisions that accumulate energy to only one of the outgoing particles, which we call energy-absorbing collisions. As we will see, this type of collisions distinguishes hard potentials from Maxwell molecules. Hence, a thorough understanding of these is crucial for the study of hard potentials.
Related literature
After the foundational works of Carleman [13, 14] ($1<p<\infty$) and Arkeryd [6] ($p=\infty$) on the propagation of $L^{p}$–norms of solutions to the homogeneous Boltzmann equation, it was Gustafsson [24, 25] who first noticed the convolution structure of the gain collisional operator by proving a Young-type inequality for $Q^{+}(f,g)$, using an intricate nonlinear interpolation procedure. Later, in a pioneering work, Lions [28] used Fourier integral operators techniques to show that $Q^{+}(f,g)$ gives a derivative gain. Wennberg [39] simplified the proof of Lions and arrived at the same estimates by using the Carleman representation and classical Fourier analysis tools. See also the works [12, 29] for another family of estimates on the quadratic operator $Q^{+}(f,f)$ that are obtained by much simpler means. Based on these results, Toscani-Villani [36] and Mouhot-Villani [32] studied the propagation of $L^{p}$–norms and smoothness of solutions to the homogeneous Boltzmann equation. However, all these early results were obtained assuming point-wise truncations in the angular variable, namely singular collisions such as grazing $(\widehat{u}\cdot\sigma=0)$ or frontal $(\widehat{u}\cdot\sigma=\pm 1)$ are discarded.
More recently, Alonso-Carneiro [1] used Fourier analysis and radial symmetrization techniques to obtain a Young-type inequality for Maxwell molecules and hard potentials, under certain integrability assumption on the cross-section $b$ that replaced point-wise truncations. The dependence on the cross-section is embedded in the constants of their estimates. Special cases of those estimates were previously obtained in [19] and [21]. Soon after that, Alonso-Carneiro-Gamba [2] extended the results of [1] to inelastic collisions, and also proved a Hardy-Littlewood-Sobolev type of inequality for soft potentials ($3<\gamma<0$).
We now focus on the most relevant prior results for our work, namely weighted convolution estimates for the gain operator without additional angular cut-offs.
Simple moment increasing estimates of this kind can be obtained by treating the operator as a Maxwell molecule with increased weights in the arguments (see [1, 2]). However, these estimates unavoidably suffer the major drawback of being moment-increasing. Namely for $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$, the estimate provided by [1, 2] is of the form:
$\|{\langle}v{\rangle}^{k}\,Q^{+}_{cl}(f,g)\|_{L^{r}_{v}}\lesssim\|{\langle}v{\rangle}^{k+\gamma}\,f\|_{L^{p}_{v}}\|{\langle}v{\rangle}^{k+\gamma}\,g\|_{L^{q}_{v}},\quad{\langle}v{\rangle}:=\sqrt{1+|v|^{2}},$
which is not sufficient to prove local/global well-posedness of the Boltzmann equation in spaces of the form ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$, or propagation of moments. The main difficulty is to offset the linear growth of the potential. Following Lions [28], a now classical (but highly technical) approach relies on decay properties of oscillatory integrals. Indeed denoting $\widehat{f}$ and $\widehat{g}$ the Fourier transforms of $f$ and $g,$ (1.9) can be written (taking $b=1$ for simplicity)
$\int_{\mathbb{R}^{9}}\widehat{f}(\eta)\widehat{g}(\xi)e^{-iV\cdot(\xi+\eta)}|u|^{\gamma}\int_{{\mathbb{S}}^{2}}e^{-i\frac{|u|}{2}(\xi-\eta)\cdot\sigma}\,d\sigma\,dv_{1}d\xi d\eta.$
Standard stationary phase considerations (see e.g. [35]) give the asymptotic behavior of the inner integral for large $|u|$:
$\int_{{\mathbb{S}}^{2}}e^{-i\frac{|u|}{2}(\xi-\eta)\cdot\sigma}\,d\sigma\approx\frac{e^{\pm i|u||\xi-\eta|}}{|u||\xi-\eta|}.$
The decay in $|u|$ compensates for the growth of the potential. The remaining challenge is to prove $L^{p}$–boundedness of the resulting Fourier integral operator. A very recent paper [26] adapts this method to the hard potential case to produce a moment-preserving estimate for the gain operator with hard potentials. However, in the case of hard-spheres, this work only gives a single estimate, namely $\|{\langle}v{\rangle}^{k}Q^{+}_{cl}(f,g)\|_{L^{6}_{v}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L^{1}_{v}}\|{\langle}v{\rangle}^{k}g\|_{L^{2}_{v}}$.
In comparison, we propose a kinetic alternative to the Fourier analytic approach, which we now present. This approach allows us to prove a much wider class of estimates.
1.2. Heuristics
For simplicity, in this subsection we focus on the case of hard-spheres i.e. $\gamma=1$ and $b=1$, which is the prototype hard potential model. We write the weighted (1.9) as
${\langle}v{\rangle}^{k}Q^{+}(f,g)(v)=2\int_{\mathbb{R}^{3}\times{\mathbb{S}}^{2}}\frac{|u|{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{k}}f_{k}(v^{*})g_{k}(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\,d\sigma dv_{1}.$
The central object in our analysis is the ratio
(1.14) $\mathcal{R}:=\frac{|u|{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{k}},\quad u=v-v_{1},\quad(v,v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}.$
We distinguish two types of collisions: energy-distributing and energy-absorbing. Each of these requires different treatment.
Energy-distributing collisions
In most instances, energy is distributed among the outgoing particles i.e. both $|v^{*}|^{2}$ and $|v_{1}^{*}|^{2}$ are comparable to the energy $E$. We call these collisions energy-distributing. Since $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$, we can then easily estimate
$\mathcal{R}\lesssim(1+E)^{\frac{1-k}{2}}\lesssim 1,$
as long as $k\geq 1$. In other words, in the most likely scenario, the weights offset the growth of the potential and the operator behaves like a Maxwell molecule. For this reason, we provide an exhaustive treatment of Maxwell molecules, which is of independent interest. In fact, we improve and extend existing results for such operators. As mentioned, convolution estimates without additional angular cut-offs for Maxwell molecules have been proved in [1, 2]. In these works, though, the implicit constants in the estimates for $b=1$ are finite only when $r<3$. We opt for a self-contained, purely kinetic approach that allows us to extend the existing estimates on Maxwell molecules for $b$ satisfying (H.1)–(H.3) to arbitrary large $r$ (even $r=\infty$), as long as $p<3$. This is the content of Theorem 3.1 ($r=\infty$) and Theorem 3.2 ($r<\infty$). We stress that our technique is also simpler, and only relies on elementary kinetic tools. Moreover, most of the aforementioned papers that introduce angular cut-offs in the gain operator fall into this category, since the cut-offs allow only energy-distributing collisions to occur. Our method could potentially be used to revisit and extend these works as well.
Energy-absorbing collisions
The second, more pathological, scenario consists of one of the outgoing particles accumulating almost all the energy of the collision. We call those energy-absorbing collisions. They distinguish hard potentials from Maxwell molecules. Following the billiard analogy, they can be visualized as two balls colliding, with one staying still after the collision while the other absorbs all of the energy.
It is easy to see that energy-absorbing collisions happen with non-trivial probability. For instance, when $\sigma\approx\frac{v}{|v|}$, the collisional law (1.3) yields $v^{*}\approx 0,\quad v_{1}^{*}\approx v$. However, in this case $\mathcal{R}\approx|u|=|v^{*}-v_{1}^{*}|\approx|v|$, where for the second equality we used (1.6). This leads to a growth of moments which cannot be controlled point-wise. Hence a crude bound would result in a moment-increasing estimate, which is not sufficient to understand the time evolution of the system, even locally in time.
However, despite the point-wise growth, we show that these singular energy-absorbing collisions do not drastically affect the averaging behavior of the gain operator. In other words, moment-preserving estimates still hold for the gain operator. The key idea is that, although energy-absorbing collisions occur with non-trivial probability, they are statistically rare. This is quantified by a point-wise cancellation and singular angular integration, see Lemma 2.5 (Cancellation Lemma) below. This estimate plays a crucial role in our analysis.
We now explain in more detail how these energy-absorbing collisions arise. They constitute the most serious obstruction to proving moment-preserving estimates.
Large velocity splitting and convolution estimates
In order to estimate $Q^{+}$, we first decompose it into two pieces $Q_{0},Q_{1}$ corresponding to the outgoing particle with the largest velocity. Namely, by conservation of energy we have either $|v^{*}|^{2}\geq E/2$ or $|v_{1}^{*}|^{2}\geq E/2$, thus we can bound
(1.15) $|Q^{+}(f,g)|\lesssim|Q_{0}(f,g)|+|Q_{1}(f,g)|,$
where
(1.16) $Q_{0}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1},$
(1.17) $Q_{1}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Therefore it suffices to prove convolution estimates for $Q_{0}$ and $Q_{1}$. This is the content of Propositions 4.14.4, which is the heart of our contribution. Upon establishing those estimates, the claim of Theorem 1.1 follows easily, see Subsection 4.5.
The main difficulty in this task is that the operators $Q_{0}$ and $Q_{1}$ are singular (in the sense that $\mathcal{R}\approx|v|$) in the regions $v_{1}^{*}\approx 0$ and $v^{*}\approx 0$ respectively, which corresponds exactly to the energy-absorbing collisions mentioned above.
As we will see, $Q_{0}$ exhibits more regular behavior due to the cut-off $\chi$ excluding some very pathological interactions. $Q_{1}$ on the other hand is more intricate since these interactions are present. Roughly speaking, $Q_{0}$ has a simple singularity, while $Q_{1}$ has a double singularity.
More specifically, the cut-off $\chi$ prevents interactions of the type $v_{1}^{*}\approx 0$ and $v_{1}\approx 0$ at the same time. To see this formally, assume $|v_{1}|\ll|v|$. Then $u=v-v_{1}\approx v$, so $|v_{1}|\ll|u|$. Using (1.3), we can write
$v_{1}^{*}=v_{1}+R_{\sigma}^{+}(u),\quad R_{\sigma}^{+}(u):=\frac{u}{2}+\frac{|u|}{2}\sigma.$
Now, the cut-off $\chi$ restricts us to the region where $\widehat{u}\cdot\sigma>0$, hence
$|R_{\sigma}^{+}(u)|^{2}=\frac{|u|^{2}}{2}\left(1+\widehat{u}\cdot\sigma\right)>\frac{|u|^{2}}{2}.$
By the triangle inequality, we obtain
$|v_{1}^{*}|\geq|R_{\sigma}^{+}(u)|-|v_{1}|>\frac{|u|}{\sqrt{2}}-|v_{1}|>\frac{|u|}{2}\gtrsim|v|,$
since $|v_{1}|\ll|u|\approx|v|$. In other words, we cannot have $v_{1}\approx 0$ and $v_{1}^{*}\approx 0$ simultaneously; therefore we can always assume that $|v_{1}|\gtrsim|v|$ when approximating the singularity of $Q_{0}$.
This reasoning does not apply to the singularity of $Q_{1}$ since
$v^{*}=v_{1}+R_{\sigma}^{-}(u),\quad R_{\sigma}^{-}(u):=\frac{u}{2}-\frac{|u|}{2}\sigma,$
and
$|R_{\sigma}^{-}(u)|^{2}=\frac{|u|^{2}}{2}\left(1-\widehat{u}\cdot\sigma\right).$
Since $\widehat{u}\cdot\sigma$ is allowed to be arbitrarily close to $1$, we can have $R_{\sigma}^{-}(u)\approx 0$, and thus $v^{*}\approx v_{1}\approx 0$. This visually corresponds to a full transfer of energy: a particle carries all the energy before the collision, it collides with a still particle, and the energy is fully transferred to the previously still particle. This situation is the most pathologic one, and results in an additional singularity which makes the estimate of $Q_{1}$ more challenging.
1.3. Main results
We now present the main results of this paper more formally. Given $k\geq 0$ and $1\leq p,q\leq\infty$, we define the normed space
(1.18) $\bm{\mathcal{X}_{p,q}^{k}}:=\left\{f\in{\langle}v{\rangle}^{-k}(L_{v}^{p}\cap L_{v}^{q})\,:\,\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}<\infty\right\},$
with norm
(1.19) $\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}:=\max\{\|{\langle}v{\rangle}^{k}f\|_{L^{p}_{v}},\|{\langle}v{\rangle}^{k}f\|_{L^{q}_{v}}\}.$
Note that the convergence of a sequence $\{f_{n}\}_{n=1}^{\infty}\subset\bm{\mathcal{X}_{p,q}^{k}}$ to some $f\in\bm{\mathcal{X}_{p,q}^{k}}$ can be characterized as follows:
[…]
It is straightforward to check that $(\bm{\mathcal{X}_{p,q}^{k}},\|\cdot\|_{\bm{\mathcal{X}_{p,q}^{k}}})$ is a Banach space.
We now state the main result of this paper:
Theorem 1.1.
Let $0<\gamma\leq 1$, $b$ satisfying (H.1)–(H.3), $1<p<2<q<\infty$, $2<r\leq\infty$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$ and $k\geq l>2$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k}L^{r}_{v}$ which satisfies the estimate
(1.21) $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},$
for all $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$.
Remark 1.2.
Unlike for polynomial weights, we do not expect moment-preserving bounds on $Q^{+}$ to hold. Indeed, consider Gaussians of the form $e^{|v|^{2}}$. To prove the corresponding weighted estimates, one would need to control the ratio
$\mathcal{R}_{exp}:=\frac{|u|^{\gamma}e^{|v|^{2}}}{e^{|v^{*}|^{2}}e^{|v_{1}^{*}|^{2}}}=|u|^{\gamma}e^{|v|^{2}-|v^{*}|^{2}-|v_{1}^{*}|^{2}}=|u|^{\gamma}e^{-|v_{1}|^{2}},$
by conservation of energy. Now for $|v|\gg 1$ and $|v_{1}|\ll 1$, we have $\mathcal{R}_{exp}\approx|v|^{\gamma}$. However, unlike for polynomial weights, this quantity does not depend on the scattering angle $\sigma.$ Therefore collisional averaging cannot offset the growth induced by the potential.
Organization of the paper
The paper is organized as follows:
  • In Section 2, we provide the necessary kinetic tools that will be used in the rest of the paper. Namely, we introduce the Bobylev variables and prove Lemma 2.2 and Lemma 2.5 which are instrumental in our analysis.
  • In Section 3, we prove convolution estimates for Maxwell molecules. Theorem 3.1 provides the convolution estimate for $r=\infty$, while Theorem 3.2 provides the convolution estimate for $r<\infty$.
  • In Section 4, we prove the convolution estimates for hard potentials, which is our most noteworthy contribution. Propositions 4.1, 4.3 provide the estimates for $Q_{0}$, while Propositions 4.2, 4.4 provide the estimates for $Q_{1}$. We conclude with the proof of Theorem 1.1.
Notation
Throughout this paper, given $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$, we will write
(1.22) $v^{*}=V-\frac{|u|}{2}\sigma,\ \ \ \ v_{1}^{*}=V+\frac{|u|}{2}\sigma,\ \ \ \ V=\frac{v+v_{1}}{2},\ \ \ u=v-v_{1},\ \ \ E=|v|^{2}+|v_{1}|^{2}.$
We will also be using the following notation:
  • We write $\chi:=\mathds{1}_{(0,+\infty)}$ for the standard cut-off function.
  • We write $C^{\infty}_{c}({\mathbb{R}}^{3})$ for the set of smooth, compactly supported functions in ${\mathbb{R}}^{3}$.
  • We write $A\lesssim B,$ to mean that there exists a numerical constant $C>0$, independent of $A,B$, such that $A\leq CB.$
  • We write $A\approx B$ to signify that both $A\lesssim B$ and $B\lesssim A$ hold.
  • We use the standard Japanese bracket notation: $\langle v\rangle:=\sqrt{1+|v|^{2}},$ where $|\cdot|$ denotes the $\ell_{v}^{2}$–norm of the vector $v\in\mathbb{R}^{3}.$
  • For $v\in\mathbb{R}^{3}$ with $v\neq 0$, we denote $\widehat{v}:=v/|v|.$
Acknowledgments
I.A. was supported by the NSF grant No. DMS-2418020 and the PSC-CUNY Research Award 68653-00 56. The authors are also grateful to the Institute for Advanced Study for providing a peaceful environment to conduct this research during July 2026.
2. Technical toolbox
In this section, we present some technical results fundamental to our analysis. First, we introduce the so-called Bobylev variables [9, 10], and then employ them to prove key angular averaging estimates. Namely, Lemma 2.2 will be used to treat Maxwell molecules, while Lemma 2.5 unveils the cancellation property of the gain collisional operator which allows us to control energy-absorbing collisions, and therefore treat hard potentials. Finally, we record the well-known pre-post collisional change of variables in Lemma 2.6.
2.1. Bobylev variables
Given $\sigma\in{\mathbb{S}}^{2}$, we define the maps $R_{\sigma}^{+},R_{\sigma}^{-}:{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ by
(2.1) $R_{\sigma}^{+}(y)$ $=\frac{y}{2}+\frac{|y|}{2}\sigma,$
(2.2) $R_{\sigma}^{-}(y)$ $=\frac{y}{2}-\frac{|y|}{2}\sigma.$
One can readily verify that the following relations hold for any $y\in{\mathbb{R}}^{3}$:
(2.3) $R_{\sigma}^{+}(y)+R_{\sigma}^{-}(y)$ $=y,$
(2.4) $R_{\sigma}^{+}(y)\cdot R_{\sigma}^{-}(y)$ $=0,$
(2.5) $|R_{\sigma}^{+}(y)|^{2}+|R_{\sigma}^{-}(y)|^{2}$ $=|y|^{2}.$
Note that given $v,v_{1}\in{\mathbb{R}}^{3}$, $\sigma\in{\mathbb{S}}^{2}$, we can connect $R_{\sigma}^{+},R_{\sigma}^{-}$ with $v^{*},v_{1}^{*}$ as follows:
(2.6) $v^{*}$ $=v-R_{\sigma}^{+}(u),$
(2.7) $v_{1}^{*}$ $=v-R_{\sigma}^{-}(u).$
We now state some finer substitution properties of $R_{\sigma}^{+},R_{\sigma}^{-}$ that will be important in uncovering the collisional averaging behavior of the gain operator. The proof of the following result can be found in e.g. the work of Arsenio [7], as well as our previous work [3]. For convenience of the reader though, we present a detailed proof.
Proposition 2.1.
Let $\sigma\in{\mathbb{S}}^{2}$ and $\epsilon\in\{+,-\}$. Then the map
$R_{\sigma}^{\epsilon}:\{y\in{\mathbb{R}}^{3}:y\cdot\sigma\neq-\epsilon|y|\}\to\{\nu\in{\mathbb{R}}^{3}:\epsilon\,(\nu\cdot\sigma)>0\},$
is a diffeomorphism with inverse
(2.8) $\big{(}R_{\sigma}^{\epsilon}\big{)}^{-1}(\nu)=2\nu-\frac{|\nu|}{(\widehat{\nu}\cdot\sigma)}\sigma,$
and Jacobian
(2.9) $\text{Jac}\,(R_{\sigma}^{\epsilon})^{-1}(\nu)=\frac{4}{(\widehat{\nu}\cdot\sigma)^{2}}.$
Moreover, for any $y\in{\mathbb{R}}^{3}$ with $y\cdot\sigma\neq-\epsilon|y|$, we have
(2.10) $|R_{\sigma}^{\epsilon}(y)\cdot\sigma|$ $=\epsilon\,(R_{\sigma}^{\epsilon}(y\cdot\sigma),$
(2.11) $|y|$ $=\frac{|R_{\sigma}^{\epsilon}(y)|}{|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|},$
(2.12) $\widehat{y}\cdot\sigma$ $=\epsilon\left(2|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|^{2}-1\right).$
Finally for $y\in{\mathbb{R}}^{3}$ with $y\cdot\sigma\neq\pm|y|$, we have
(2.13) $|\widehat{R}_{\sigma}^{+}(y)|^{2}+|\widehat{R}_{\sigma}^{-}(y)|^{2}=1.$
Proof.
Let $\sigma\in{\mathbb{S}}^{2}$ and $\epsilon\in\{+,-\}$. Let us also denote
$A_{\sigma}^{\epsilon}$ $=\{y\in{\mathbb{R}}^{3},\>\,y\cdot\sigma\neq-\epsilon|y|\},$
$B_{\sigma}^{\epsilon}$ $=\{\nu\in{\mathbb{R}}^{3}\,:\,\epsilon\,(\nu\cdot\sigma)>0\}.$
First, for $y\in A_{\sigma}^{\epsilon}$ we clearly have $R_{\sigma}^{\epsilon}(y)\cdot\sigma\neq 0$. Moreover, since $\epsilon\in\{+,-\}$, we obtain
(2.14) $\begin{split}\epsilon(R_{\sigma}^{\epsilon}(y)\cdot\sigma)&=\epsilon\left(\frac{y\cdot\sigma}{2}+\epsilon\frac{|y|}{2}\right)=\frac{|y|}{2}(1+\epsilon\,\widehat{y}\cdot\sigma)\\ &=\frac{|y|}{2}|1+\epsilon\,\widehat{y}\cdot\sigma|=\frac{|y|}{2}|\epsilon+\widehat{y}\cdot\sigma|\\ &=\left|\frac{y\cdot\sigma}{2}+\epsilon\frac{|y|}{2}\right|=|R_{\sigma}^{\epsilon}(y)\cdot\sigma|,\end{split}$
thus $R_{\sigma}^{\epsilon}:A_{\sigma}^{\epsilon}\to B_{\sigma}^{\epsilon}$, as well as (2.10) holds.
In order to show that $R_{\sigma}^{\epsilon}$ is a diffeomorphism with inverse given by (2.8), consider $y\in A_{\sigma}^{\epsilon}$, $\nu\in B_{\sigma}^{\epsilon}$ with $\nu=R_{\sigma}^{\epsilon}(y)$. Then $y=2\nu-\epsilon|y|\sigma$, thus
$|y|^{2}=4|\nu|^{2}-4\epsilon|y|(\nu\cdot\sigma)+|y|^{2}\Rightarrow|y|=\frac{|\nu|}{\epsilon(\widehat{\nu}\cdot\sigma)},$
and (2.8) follows. Moreover, it is clear that $R_{\sigma}^{\epsilon},(R_{\sigma}^{\epsilon})^{-1}$ are continuously differentiable in $A_{\sigma}^{\epsilon},B_{\sigma}^{\epsilon}$ respectively, thus (2.8) follows. Moreover, it is clear that $R_{\sigma}^{\epsilon},\,(R_{\sigma}^{\epsilon})^{-1}$ are continuously differentiable in $A_{\sigma}^{\epsilon},B_{\sigma}^{\epsilon}$, thus $R_{\sigma}^{\epsilon}:A_{\sigma}^{\epsilon}\to B_{\sigma}^{\epsilon}$ is a diffeomorphism.
To compute the Jacobian, we differentiate (2.8) in $\nu$, to obtain
$D(R_{\sigma})^{-1}(\nu)=2I_{3}-\sigma\nabla^{T}f(\nu),$
where $I_{3}$ is the identity matrix and $f(\nu)=\frac{|\nu|^{2}}{(\nu\cdot\sigma)}$. Thus, we have111we use the Linear Algebra identity $\det(\lambda I_{n}+vw^{T})=\lambda^{n}(1+\lambda^{-1}v\cdot w)$, where $v,w\in{\mathbb{R}}^{n}$, $n\in\mathbb{N}$ and $\lambda\neq 0$.
(2.15) $\text{Jac}(R_{\sigma})^{-1}(\nu)=\det\left(2I_{3}-\sigma\nabla^{T}f(\nu)\right)=8\left(1-\frac{1}{2}\nabla f(\nu)\cdot\sigma\right).$
We readily compute
$\nabla f(\nu)\cdot\sigma=\frac{2(\nu\cdot\sigma)^{2}-|\nu|^{2}}{(\nu\cdot\sigma)^{2}}=2-\frac{|\nu|^{2}}{(\nu\cdot\sigma)^{2}},$
which, combined with (2.15), gives (2.9).
It remains to prove (2.11)–(2.13). Indeed, since $y=2R_{\sigma}^{\epsilon}(y)-\epsilon|y|\sigma$, we have
$|y|^{2}=\Big{|}2R_{\sigma}^{\epsilon}(y)-\epsilon|y|\sigma\Big{|}^{2}=4|R_{\sigma}^{\epsilon}(y)|^{2}+|y|^{2}-4\epsilon|y|(R_{\sigma}^{\epsilon}(y)\cdot\sigma),$
thus using (2.10) we obtain
$|R_{\sigma}^{\epsilon}(y)|^{2}=\epsilon|y|(R_{\sigma}^{\epsilon}(y)\cdot\sigma)=|y|\,|R_{\sigma}^{\epsilon}(y)\cdot\sigma|=|y|\,|R_{\sigma}^{\epsilon}(y)|\,|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|,$
and (2.11) follows.
Now using (2.11), (2.10), we obtain
$y\cdot\sigma$ $=2(R_{\sigma}^{\epsilon}(y)\cdot\sigma)-\epsilon|y|=\epsilon|y|\left(2\frac{\epsilon(R_{\sigma}^{\epsilon}(y)\cdot\sigma)}{|y|}-1\right)$
$=\epsilon|u|\left(2\frac{|R_{\sigma}^{\epsilon}(y)\cdot\sigma|}{|y|}-1\right)=\epsilon|y|\left(2|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|^{2}-1\right),$
and (2.12) follows.
Finally, for $y\in{\mathbb{R}}^{3}$ with $y\cdot\sigma\neq\pm|y|$ we obtain (2.13) by combining (2.5) with (2.11). ∎
2.2. Maxwell collisional averaging
Here, we use Lemma 2.1 to obtain the main averaging estimates for Maxwell molecules:
Lemma 2.2.
Let $\beta_{1}<1$, $\beta_{2}>1/2$ and $1\leq p<\infty$. For $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$, we denote
(2.16) $\mathcal{I}^{1}_{\beta_{1},p}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{1}}}\,d\sigma\,dv_{1},$
(2.17) $\mathcal{I}^{2}_{\beta_{2},p}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v_{1}^{*})|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{2}}\,d\sigma\,dv_{1}.$
Then, there hold the estimates
(2.18) $\|\mathcal{I}_{\beta_{1},p}^{1}[f]\|_{L^{\infty}_{v}}$ $\lesssim\|f\|_{L^{p}_{v}}^{p},$
(2.19) $\|\mathcal{I}_{\beta_{2},p}^{2}[f]\|_{L^{\infty}_{v}}$ $\lesssim\|f\|_{L^{p}_{v}}^{p}.$
Proof.
Consider $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$. We first prove (2.18). Fix $v\in{\mathbb{R}}^{3}$. Using (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ we write
$\mathcal{I}_{\beta_{1},p}^{1}[f](v)$ $=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(u))|^{p}\chi(2|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{\beta_{1}}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(y))|^{p}\chi(2|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{\beta_{1}}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ and obtain
$\mathcal{I}_{\beta_{1},p}^{1}[f](v)$ $\approx\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-\nu)|^{p}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{\beta_{1}}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\int_{{\mathbb{S}}^{2}}\frac{\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{\beta_{1}}}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{0}^{1}\frac{\chi(2x^{2}-1)}{x^{2}(1-x^{2})^{\beta_{1}}}\,dx\right)\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{1/\sqrt{2}}^{1}\frac{1}{(1-x^{2})^{\beta_{1}}}\,dx\right)\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{1/\sqrt{2}}^{1}\frac{1}{(1-x)^{\beta_{1}}}\,dx\right)\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\,d\nu=\|f\|_{L^{p}_{v}}^{p},$
where we used the fact that $\beta_{1}<1$ for the convergence of the integral in $x$ and translation invariance for the last equality. Estimate (2.18) is proved.
We now prove (2.26). Fix again $v\in{\mathbb{R}}^{3}$. Using (2.7) and (2.12), we write
$\mathcal{I}_{\beta_{2},p}^{2}[f](v)$ $=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f(v-R_{\sigma}^{-}(u))|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{2}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f(v-R_{\sigma}^{-}(y))|^{p}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|^{2\beta_{2}}\,dy\,d\sigma.$
Using again Proposition 2.1 to substitute $\nu:=R_{\sigma}^{-}(y)$, we obtain
$\mathcal{I}_{\beta_{2},p}^{2}[f](v)$ $\approx\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}|\widehat{\nu}\cdot\sigma|^{2(\beta_{2}-1)}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\int_{{\mathbb{S}}^{2}}|\widehat{\nu}\cdot\sigma|^{2(\beta_{2}-1)}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{0}^{1}x^{2(\beta_{2}-1)}\,dx\right)\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\,d\nu=\|f\|_{L^{p}_{v}}^{p},$
where we used the fact that $\beta_{2}>1/2$ for the convergence of the integral in $x$. Estimate (2.19) is proved. ∎
2.3. Cancellation Lemma
Now, we use Lemma 2.1 to prove Lemma 2.5 which encodes the cancellation mechanism of the gain collisional operator. First, we present an important point-wise cancellation estimate in Lemma 2.3 and a singular integral estimate in Lemma 2.4:
Lemma 2.3.
For any $v,v_{1}\in{\mathbb{R}}^{3}$, $\sigma\in{\mathbb{S}}^{2}$, the following bounds holds
(2.20) ${\langle}v^{*}{\rangle}$ $\geq\frac{{\langle}v_{1}^{*}{\rangle}}{\sqrt{2}}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|^{2}\right)^{1/2},\quad\lambda_{v_{1}^{*}}=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}},$
(2.21) ${\langle}v_{1}^{*}{\rangle}$ $\geq\frac{{\langle}v^{*}{\rangle}}{\sqrt{2}}\left(1-\lambda_{v^{*}}|\widehat{v^{*}}\cdot\sigma|^{2}\right)^{1/2},\quad\lambda_{v^{*}}=\frac{|v^{*}|^{2}}{{\langle}v^{*}{\rangle}^{2}}.$
Proof.
Fix $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$. Using the elementary inequality
$|u||V|\leq\frac{|u|^{2}}{4}+|V|^{2}=\frac{E}{2},$
we obtain
$|v^{*}|^{2}$ $=\frac{E}{2}-|u||V|(\widehat{V}\cdot\sigma)\geq\frac{E}{2}(1-|\widehat{V}\cdot\sigma|)=\frac{E}{2}\frac{1-|\widehat{V}\cdot\sigma|^{2}}{1+|\widehat{V}\cdot\sigma|}\geq\frac{E}{4}(1-|\widehat{V}\cdot\sigma|^{2})$
$\geq\frac{|V|^{2}}{2}(1-|\widehat{V}\cdot\sigma|^{2}).$
Now, using the expression $V=v_{1}^{*}-\frac{|u|}{2}\sigma$, we compute
$|V|^{2}(1-|\widehat{V}\cdot\sigma|^{2})$ $=\left|v_{1}^{*}-\frac{|u|}{2}\sigma\right|^{2}-\left|v_{1}^{*}\cdot\sigma-\frac{|u|}{2}\right|^{2}=|v_{1}^{*}|^{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|^{2}),$
thus $|v^{*}|^{2}\geq\frac{|v_{1}^{*}|^{2}}{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|^{2})$. Hence, we obtain
${\langle}v_{1}^{*}{\rangle}^{2}$ $=1+|v_{1}^{*}|^{2}\geq\frac{1}{2}\left(1+|v_{1}^{*}|^{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|^{2})\right)=\frac{{\langle}v_{1}^{*}{\rangle}^{2}}{2}(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|^{2}),$
where $\lambda=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}}$, and (2.20) is proved.
The bound (2.21) is proved identically using the expression $V=v^{*}+\frac{|u|}{2}\sigma$ instead. ∎
Lemma 2.4.
Let $l>2$. Then for any $\widehat{\eta}\in{\mathbb{S}}^{2}$ and any $0<\lambda<1$, the following bound holds
(2.22) $\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|^{2})^{l/2}}\,d\sigma\lesssim\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2}.$
Proof.
Let $l>2$ and fix $\eta\in{\mathbb{S}}^{2}$ as well as $0<\lambda<1$. Integrating in spherical coordinates, we obtain
$\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|^{2})^{l/2}}\,d\sigma$ $\approx\int_{0}^{1}\frac{1}{(1-\lambda x^{2})^{l/2}}\,dx\leq\int_{0}^{1}\frac{1}{(1-\lambda x)^{l/2}}\,dx$
$=\frac{1}{\lambda}\int_{0}^{\lambda}\frac{1}{(1-z)^{l/2}}\,dz\leq\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2},$
since $l>2$. ∎
We are now in the position to prove Lemma 2.5:
Lemma 2.5 (Cancellation lemma).
Let $l>2$ and $1\leq p<\infty$. For $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$, we denote
(2.23) $\mathcal{J}_{l,p}^{1}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>4}\mathds{1}_{|v^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1},$
(2.24) $\mathcal{J}_{l,p}^{2}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)|^{2}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Then there holds the estimates
(2.25) $\|{\langle}v{\rangle}^{2}\mathcal{J}_{l,p}^{1}[f]\|_{L_{v}^{\infty}}$ $\lesssim\|f\|_{L_{v}^{p}}^{p},$
(2.26) $\|{\langle}v{\rangle}^{2}\mathcal{J}_{l,p}^{2}[f]\|_{L_{v}^{\infty}}$ $\lesssim\|f\|_{L_{v}^{p}}^{p}.$
Proof.
Throughout the proof, given $a\in{\mathbb{R}}^{3}$ with $a\neq 0$, we will denote $\lambda_{a}:=\frac{|a|^{2}}{{\langle}a{\rangle}^{2}}\in(0,1)$. Consider $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$.
Estimate for $\mathcal{J}_{l,p}^{1}$
We first prove (2.25). Fix $v\in{\mathbb{R}}^{3}$. We use (2.21), followed by (2.6) and (2.12) to bound
$\mathcal{J}_{l,p}^{1}[f](v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{{\langle}v^{*}{\rangle}^{l}\left(1-\lambda_{v^{*}}|\widehat{v^{*}}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v^{*}|>|v|/2>2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(u))|^{p}\chi(2|\widehat{R}^{+}_{\sigma}(u)\cdot\sigma|^{2}-1)}{\left\langle v-R_{\sigma}^{+}(u)\right\rangle^{l}\left(1-\lambda_{v-R_{\sigma}^{+}(u)}|\widehat{(v-R_{\sigma}^{+}(u))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{+}(u)|>|v|/2>2}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(y))|^{p}\chi(2|\widehat{R}^{+}_{\sigma}(y)\cdot\sigma|^{2}-1)}{{\langle}v-R_{\sigma}^{+}(y){\rangle}^{l}\left(1-\lambda_{v-R_{\sigma}^{+}(y)}|\widehat{(v-R_{\sigma}^{+}(y))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{+}(y)|>|v|/2>2}\,dy\,d\sigma.$
Using Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$, we obtain
$\mathcal{J}_{l,p}^{1}[f](v)\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{-l}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\frac{\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l/2}}\,d\nu\,d\sigma$
$\approx\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{-l}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\int_{{\mathbb{S}}^{2}}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l/2}}\,d\sigma\,d\nu,$
where we used the basic inequality $\frac{\chi(2z^{2}-1)}{z^{2}}\lesssim 1$, $z\in[0,1]$.
Then, since $l>2$ and $\lambda_{v-\nu}=\frac{|v-\nu|^{2}}{{\langle}v-\nu{\rangle}^{2}}\in(0,1)$, Lemma 2.4 implies
$\mathcal{J}_{l,p}^{1}[g](v)$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{-l}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\frac{(1-\lambda_{v-\nu})^{1-l/2}}{\lambda_{v-\nu}}\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}\frac{|f(v-\nu)|^{p}}{|v-\nu|^{2}}\mathds{1}_{|v-\nu|>|v|/2>2}\,d\nu$
$\approx{\langle}v{\rangle}^{-2}\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\,d\nu={\langle}v{\rangle}^{-2}\|f\|_{L_{v}^{p}}^{p}.$
Estimate (2.26) is proved.
Estimate for $\mathcal{J}_{l,p}^{2}$
Fix $v\in{\mathbb{R}}^{3}$. We use (2.20) and (2.7) to bound
$\mathcal{J}_{l,p}^{2}[f](v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v_{1}^{*})|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v_{1}^{*}|>|v|/2>2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{-}(u))|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{\left\langle v-R_{\sigma}^{-}(u)\right\rangle^{l}\left(1-\lambda_{v-R_{\sigma}^{+}(u)}|\widehat{(v-R_{\sigma}^{-}(u))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(u)|>|v|/2>2}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{-}(y))|^{p}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|^{2}}{{\langle}v-R_{\sigma}^{-}(y){\rangle}^{l}\left(1-\lambda_{v-R_{\sigma}^{-}(y)}|\widehat{(v-R_{\sigma}^{-}(y))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(y)|>|v|/2>2}\,dy\,d\sigma.$
Using Proposition 2.1 to substitute $\nu:=R_{\sigma}^{-}(y)$, we obtain
$\mathcal{J}_{l,p}^{2}[f](v)\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{\alpha}{\langle}v-\nu{\rangle}^{-l\alpha}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l\alpha/2}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{\alpha}{\langle}v-\nu{\rangle}^{-l\alpha}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\int_{{\mathbb{S}}^{2}}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l\alpha/2}}\,d\sigma\,d\nu.$
Following an identical argument to the one used for $\mathcal{J}_{l,p}^{1}$, we conclude (2.26). ∎
2.4. Pre-post collisional change of variables
Finally, we record the well-known pre-post collisional change of variables, see e.g. [37], which will be useful in the $L^{r}$– estimates. We provide the proof for convenience of the reader.
Lemma 2.6.
The map $T:(v,v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\to(v^{*},v_{1}^{*},\eta)\in{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}$ where
(2.27) $\begin{cases}v^{*}=\frac{v+v_{1}}{2}-\frac{|v-v_{1}|}{2}\sigma\\ v_{1}^{*}=\frac{v+v_{1}}{2}+\frac{|v-v_{1}|}{2}\sigma\\ \eta=\frac{v_{1}-v}{|v-v_{1}|}\end{cases}$
is an involution of ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}$. Moreover, for any non-negative and continuously differentiable function $F:{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times[-1,1]\to{\mathbb{R}}_{+}$, there holds the change of variables formula
(2.28) $\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v^{*},v_{1}^{*},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v,v_{1},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv.$
Remark 2.7.
It is customary in literature to abbreviate notation and write $(v,v_{1})\to(v^{*},v_{1}^{*})$ to indicate the above change of variables, however we stress that it is not correct to perform this change of variables for a fixed $\sigma\in{\mathbb{S}}^{2}$.
Proof.
Writing $T^{2}:=T\circ T$, we compute
$T^{2}(v,v_{1},\sigma)$ $=T(v^{*},v_{1}^{*},\eta)=\begin{pmatrix}\frac{v^{*}+v_{1}^{*}}{2}-\frac{|v^{*}-v_{1}^{*}|}{2}\eta\\ \frac{v^{*}+v_{1}^{*}}{2}+\frac{|v^{*}-v_{1}^{*}|}{2}\eta\\ \frac{v_{1}^{*}-v^{*}}{|v^{*}-v_{1}^{*}|}\end{pmatrix}=\begin{pmatrix}\frac{v+v_{1}}{2}-\frac{|v-v_{1}|}{2}\eta\\ \frac{v+v_{1}}{2}+\frac{|v-v_{1}|}{2}\eta\\ \frac{v_{1}^{*}-v^{*}}{|v-v_{1}|}\end{pmatrix}=\begin{pmatrix}v\\ v_{1}\\ \sigma\end{pmatrix}$.
Therefore $T^{2}=Id$, so $T$ is an involution.
Second, writing $u^{*}:=v^{*}-v_{1}^{*}$, we compute
$u^{*}\cdot\eta=-|v-v_{1}|\frac{\sigma\cdot(v_{1}-v)}{|v-v_{1}|}=u\cdot\sigma,$
thus by (1.6) we obtain $\widehat{u}\cdot\sigma=\widehat{u^{*}}\cdot\eta.$ Therefore
$\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v^{*},v_{1}^{*},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v^{*},v_{1}^{*},\widehat{u^{*}}\cdot\eta)\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v,v_{1},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv,$
since $T$ is an involution and thus has unitary Jacobian. ∎
3. Convolution estimates for Maxwell molecules
In the section, we prove the convolution estimates for Maxwell molecules. Given $b$ satisfying (H.1)–(H.3), we define the Maxwell collisional operator with angular cross-section $b$ as
(3.1) $Q_{M,b}^{+}(f,g)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Slightly abbreviating notation we will be writing $Q^{+}_{M}:=Q^{+}_{M,1}$ for the Maxwell operator with constant cross-section equal to $1$.
As mentioned, when $b$ is constant, the previous works [1, 2] provide an estimate only when $r<3$. Here, we are able to extend the estimates to arbitrary $r$, as long as $p<3$.
3.1. The $L_{v}^{\infty}$– estimate for Maxwell molecules
First, we prove the convolution estimate corresponding to $r=\infty$.
Theorem 3.1.
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{\infty}_{v}$ satisfying the estimate
(3.2) $\|Q^{+}_{M,b}\|_{L_{v}^{\infty}}\lesssim\|f\|_{L_{v}^{p}}\|g\|_{L_{v}^{q}}.$
For $p=1$, $q=\infty$, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{1}_{v}\times C_{b}^{1}({\mathbb{R}}^{3})\to L^{\infty}_{v}$ satisfying the estimate
(3.3) $\|Q^{+}_{M,b}\|_{L^{\infty}_{v}}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}},$
where $C_{b}^{1}({\mathbb{R}}^{3})$ denotes the space of continuously differentiable and bounded functions in $v$.
Proof.
We start by proving the first claim. By density, it suffices to show (3.2) for $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Since $p<3$, we have $\frac{1}{2q}<\frac{1}{p}$. Pick $\alpha>1/2$ with $\alpha p<q$. We use (H.3) and Hölder’s inequality to bound
$|Q_{M,b}^{+}(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f(v^{*})|\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/q}}\right)\left(|g(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/q}\right)\,d\sigma\,dv_{1}$
(3.4) $\leq I_{1}^{1/p}(v)I_{2}^{1/q}(v),$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha p/q}}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g(v_{1}^{*})|^{q}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}\,d\sigma\,dv_{1}.$
Since $\alpha p<q$, $\alpha>1/2$, Lemma 2.2 implies $I_{1}(v)\lesssim\|f\|_{L^{p}_{v}}^{p}$ and $I_{2}(v)\lesssim\|g\|_{L_{v}^{q}}^{q}$. Combining (3.4) with the bounds for $I_{1},I_{2}$, (3.2) follows. The claim is proved.
Now for the case $p=1$, $q=\infty$, again by density, it suffices to show (3.3) for $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Assume $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$ and $g\in C_{b}^{1}({\mathbb{R}}^{3})$. By (H.3) and (2.18), we bound
(3.5) $|Q^{+}_{M,b}(f,g)|$ $\leq\|g\|_{L^{\infty}_{v}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}},$
and (3.3) follows. The claim is proved.
3.2. The $L_{v}^{r}$– estimate for Maxwell molecules
Now, we prove the convolution estimate for general $1\leq r<\infty$.
Theorem 3.2.
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q,r<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{r}_{v}$ satisfying the estimate
(3.6) $\|Q_{M,b}^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f\|_{L_{v}^{p}}\|g\|_{L_{v}^{q}}.$
Proof.
By density, it suffices to prove (3.6) for $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. We take different cases for $1<p<3$ and $p=1$.
$\bullet$ Case $1<p<3$ : Since $p<3$, we have $p^{\prime}>3/2$, thus
(3.7) $\frac{1}{2p^{\prime}}=\frac{1}{2}-\frac{1}{2q^{\prime}}-\frac{1}{2r}=\frac{1}{q^{\prime}}+\frac{1}{r}+\frac{1}{2}-\frac{3}{2q^{\prime}}-\frac{3}{2r}=\frac{1}{q^{\prime}}+\frac{1}{r}+\frac{3}{2p^{\prime}}-1<\frac{1}{q^{\prime}}+\frac{1}{r}.$
Take $\alpha>1/2$ with $\frac{\alpha}{p^{\prime}}<\frac{1}{q^{\prime}}+\frac{1}{r}$, which is possible due to the above bound. Equivalently, we can write this relation as $1-\frac{p^{\prime}}{\alpha q^{\prime}}<\frac{p^{\prime}}{\alpha r}$. Pick $s\in{\mathbb{R}}$ with $1-\frac{p^{\prime}}{\alpha q^{\prime}}<s<\frac{p^{\prime}}{\alpha r}$. Then $\frac{(1-s)\alpha q^{\prime}}{p^{\prime}}<1$ and $\frac{s\alpha r}{p^{\prime}}<1$.
Fix $v\in{\mathbb{R}}^{3}$. Using (H.3) and Hölder’s inequality, we bound
$|Q_{M,b}^{+}(f,g)(v)|$ $\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f(v^{*})|^{p/q^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha(1-s)/p^{\prime}}}\right)\left(|g(v_{1}^{*})|^{q/p^{\prime}}|\widehat{R}^{-}_{\sigma}(u)\cdot\sigma|^{2\alpha/p^{\prime}}\right)$
$\hskip 113.81102pt\times\left(\frac{|f(v^{*})|^{p/r}|g(v_{1}^{*})|^{q/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha s/p^{\prime}}}\right)\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(3.8) $\lesssim I_{1}^{1/q^{\prime}}(v)I_{2}^{1/p^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|g(v_{1}^{*})|^{q}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha sr/p^{\prime}}}\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha(1-s)q^{\prime}/p^{\prime}}}\,d\sigma\,dv_{1}\,dv,$
$I_{2}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g(v_{1}^{*})|^{q}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}\,d\sigma\,dv_{1}\,dv.$
Since $\frac{(1-s)\alpha q^{\prime}}{p^{\prime}}<1$ and $\alpha>1/2$, Lemma 2.2 implies $I_{1}(v)\lesssim\|f\|_{L^{p}_{v}}^{p}$ and $I_{2}(v)\lesssim\|g\|_{L^{q}}^{q}$.
Combining (3.8) with the bounds for $I_{1},I_{2}$, we obtain
$|Q_{M,b}^{+}(f,g)(v)|\lesssim\|f\|_{L_{v}^{p}}^{p/q^{\prime}}\|g\|_{L_{v}^{q}}^{q/p^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|g(v_{1}^{*})|^{q}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha sr/p^{\prime}}}\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above bound to the $r$ and integrating in $v$, and using (2.28), we obtain
$\|Q_{M,b}^{+}(f,g)\|_{L_{v}^{r}}^{r}$ $\lesssim\|f\|_{L_{v}^{p}}^{r-p}\|g\|_{L_{v}^{q}}^{r-q}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|g(v_{1}^{*})|^{q}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha sr/p^{\prime}}}\,d\sigma\,dv_{1}\,dv$
$=\|f\|_{L_{v}^{p}}^{r-p}\|g\|_{L_{v}^{q}}^{r-q}\int_{{\mathbb{R}}^{6}}|f(v)|^{p}|g(v_{1})|^{q}\left(\int_{{\mathbb{S}}^{2}}\frac{\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)|^{2\alpha sr/p^{\prime}}}\,d\sigma\right)\,dv_{1}\,dv.$
To estimate the inner sphere integral we use (2.12) to write
$\int_{{\mathbb{S}}^{2}}\frac{\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)|^{2\alpha sr/p^{\prime}}}\,d\sigma\approx\int_{{\mathbb{S}}^{2}}\frac{\chi(\widehat{u}\cdot\sigma)}{(1-\widehat{u}\cdot\sigma)^{s\alpha r/p^{\prime}}}\,d\sigma\approx\int_{0}^{1}\frac{1}{(1-x)^{s\alpha r/p^{\prime}}}\,dx\approx 1,$
since $\frac{s\alpha r}{p^{\prime}}<1$.
We conclude
(3.9) $\|Q_{M,b}^{+}(f,g)\|_{L^{r}_{v}}^{r}\lesssim\|f\|_{L_{v}^{p}}^{r-p}\|g\|_{L_{v}^{q}}^{r-q}\int_{{\mathbb{R}}^{6}}|f(v)|^{p}|g(v_{1})|^{q}\,dv_{1}\,dv=\|f\|_{L_{v}^{p}}^{r}\|g\|_{L_{v}^{q}}^{r}.$
$\bullet$ Case $p=1$ : We use (H.3) and Hölder’s inequality to bound
$|Q_{M,b}^{+}(f,g)(v)|$ $\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|^{1/r^{\prime}}\left(|f(v^{*})|^{1/r}|g(v_{1}^{*})|\right)\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(3.10) $\leq I_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
By (2.18), we have $I_{1}\lesssim\|f\|_{L^{1}_{v}}$. Combining (3.10) with the estimate on $I_{1}$, we obtain
$|Q_{M,b}^{+}(f,g)(v)|\lesssim\|f\|_{L^{1}_{v}}^{1/r^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above expression to the $r$-th power and integrating in $v$, we obtain
(3.11) $\|Q_{M,b}^{+}(f,g)(v)\|_{L^{r}_{v}}^{r}$ $\lesssim\|f\|_{L^{1}_{v}}^{r-1}\left(\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\,dv\right)\lesssim\|f\|_{L^{1}_{v}}^{r}\|g\|_{L^{r}_{v}}^{r}.$
Combining (3.9), (3.11), estimate (3.6) follows. The claim is proved. ∎
4. Convolution estimates for hard-spheres
In this final section, we prove our main result, Theorem 1.1. As mentioned in Subsection 1.2, we decompose $Q^{+}$ into two pieces $Q_{0},\,Q_{1}$ corresponding to the outgoing particle with the largest velocity, and estimate each piece separately. This is the content of Propositions 4.14.4 which are the heart of our contributions. Upon establishing those estimates, the claim of Theorem 1.1 follows easily.
Namely, given $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$, conservation of energy implies that either $|v^{*}|^{2}\geq E/2$ or $|v_{1}^{*}|^{2}\geq E/2$. Thus, using (H.3), we can bound
(4.1) $|Q^{+}(f,g)|\lesssim|Q_{0}(f,g)|+|Q_{1}(f,g)|,$
where
(4.2) $Q_{0}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1},$
(4.3) $Q_{1}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
In the following, we prove convolution estimates for the operators $Q_{0},Q_{1}$.
4.1. The $L_{v}^{\infty}$– estimate for $Q_{0}$
Proposition 4.1.
Let $0<\gamma\leq 1$. Consider $1\leq p_{0},q_{0}\leq\infty$ with $p_{0}\leq 2$, such that $\frac{1}{p_{0}}+\frac{1}{q_{0}}=1$. Then for any $k\geq l>2$, there holds the estimate
(4.4) $\|{\langle}v{\rangle}^{k}Q_{0}(f,g)\|_{L_{v}^{\infty}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p_{0}}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q_{0}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{k}:={\langle}v{\rangle}^{k}f$, $g_{l}:={\langle}v{\rangle}^{l}g$. By the triangle inequality, we can bound $Q_{0}$ as follows
(4.5) $|Q_{0}(f,g)|\lesssim|Q_{0}^{sm}(f,g)|+|Q_{0}^{r}(f,g)|+|Q_{0}^{s}(f,g)|,$
where
$Q_{0}^{sm}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|\leq 10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{0}^{r}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|v_{1}^{*}|\geq|v|/10}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{0}^{s}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Estimate for $Q_{0}^{sm}$
Using the inequality $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we bound
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|\leq 10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\lesssim E^{1/2}(1+E)^{-k/2}\lesssim 1,$
since $k>2$. Using the above estimate, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{sm}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})g_{l}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}=Q_{M}^{+}(|f_{k}|,|g_{l}|),$
where we recall the Maxwell operator notation from (3.1). Then, Theorem 3.1 implies
(4.6) $\|{\langle}v{\rangle}^{k}Q_{0}^{sm}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},$
since $p_{0}\leq 2$.
Estimate for $Q_{0}^{r}$
Using the inequality $|u|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we bound
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|\geq|v|/10}\mathds{1}_{|v^{*}|>E/2}\lesssim{\langle}v{\rangle}^{k-l}E^{1/2}(1+E)^{-k/2}\lesssim(1+E)^{\frac{1-l}{2}},$
since $k\geq l>2$. Arguing as in the previous case, we obtain
(4.7) $\|{\langle}v{\rangle}^{k}Q_{0}^{r}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},$
since $p_{0}\leq 2$.
We now estimate the singular operator $Q_{0}^{s}$.
Estimate for $Q_{0}^{s}$
Using the inequality $|u|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we estimate
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\lesssim\frac{{\langle}v{\rangle}^{3}}{{\langle}v^{*}{\rangle}^{2}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}$
$\lesssim\frac{{\langle}v{\rangle}^{3}}{{\langle}u{\rangle}^{2}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\leq\frac{{\langle}v{\rangle}^{3}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}R_{\sigma}^{-}(u){\rangle}^{2}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10},$
where for the last inequality we used (2.11). However, when $|v_{1}^{*}|<|v|/10$, (2.7) and the triangle inequality imply
$|R_{\sigma}^{-}(u)|=|v-v_{1}^{*}|\geq|v|-|v_{1}^{*}|\geq\frac{9|v|}{10},$
therefore
(4.8) $\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\lesssim\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}.$
We take different cases for $1<p_{0}\leq 2$ and $p_{0}=1$.
$\bullet$ Case $1<p_{0}\leq 2$ : Using (4.8) and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})g_{l}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}|f_{k}(v^{*})|}{{\langle}v_{1}^{*}{\rangle}^{l}}\right)\left(|g_{l}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}\right)\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.9) $\leq I_{1}^{1/p_{0}}(v)I_{2}^{1/q_{0}}(v),$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{p_{0}}|f_{k}(v^{*})|^{p_{0}}}{{\langle}v_{1}^{*}{\rangle}^{lp_{0}}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g_{l}(v_{1}^{*})|^{q_{0}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q_{0}}\,d\sigma\,dv_{1}.$
Since $l>2$ and $p_{0}\leq 2$, (2.25) implies $I_{1}(v)\lesssim{\langle}v{\rangle}^{p_{0}-2}\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}\leq\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}$. By (2.19), we also have $I_{2}(v)\lesssim\|g_{l}\|_{L^{q_{0}}}^{q_{0}}$.
Using (4.9) and the estimates for $I_{1},I_{2}$, we obtain
(4.10) ${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},\quad 1<p_{0}\leq 2.$
$\bullet$ Case $p_{0}=1$ : We use again (4.8) to bound
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|$ $\lesssim\|g_{l}\|_{L_{v}^{\infty}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|f_{k}(v^{*})|}{{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.11) $\lesssim\|f_{k}\|_{L_{v}^{1}}\|g_{l}\|_{L_{v}^{\infty}},$
where for the last inequality we used again the fact that $l>2$ and (2.25).
By (4.10), (4.11), we conclude
(4.12) $\|{\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},\quad 1\leq p_{0}\leq 2.$
Conclusion of the proof
Combining (4.6), (4.7), (4.12), estimate (4.4) follows. ∎
4.2. The $L_{v}^{\infty}$– estimate for $Q_{1}$
Proposition 4.2.
Let $0<\gamma\leq 1$. Consider $1<p_{1},q_{1}<\infty$ with $p_{1}<2$ and $1\leq p_{2},q_{2}\leq\infty$ with $p_{2}<2$, such that $\frac{1}{p_{1}}+\frac{1}{q_{1}}=\frac{1}{p_{2}}+\frac{1}{q_{2}}=1$. Then for any $k\geq l>2$, there holds the estimate
(4.13) $\|{\langle}v{\rangle}^{k}Q_{1}(f,g)\|_{L_{v}^{\infty}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p_{1}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q_{1}}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q_{2}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p_{2}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$. We note that since $p_{1}<2$ and $p_{2}<2$, we have $p_{1}<2<q_{1}$ and $p_{2}<2<q_{2}$ respectively.
We use the triangle inequality to bound $Q_{1}$ as follows
(4.14) $|Q_{1}(f,g)|\lesssim|Q_{1}^{sm}(f,g)|+|Q_{1}^{r}(f,g)|+|Q_{1}^{s_{1}}(f,g)|+|Q_{1}^{s_{2}}(f,g)|,$
where
$Q_{1}^{sm}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|\leq 10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{1}^{r}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|v^{*}|\geq|v_{1}^{*}|/10}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{1}^{s_{1}}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{1}^{s_{2}}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
and given $v\in{\mathbb{R}}^{3}$, we define the singular sets
(4.15) $\mathcal{B}_{1}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v_{1}^{*}|}{10},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|\geq{\langle}v{\rangle}^{-1/2}\right\},$
(4.16) $\mathcal{B}_{2}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v_{1}^{*}|}{10},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<{\langle}v{\rangle}^{-1/2}\right\}.$
Estimate for $Q_{1}^{sm}$
Using the inequality $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we bound
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v|\leq 10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\lesssim E^{1/2}(1+E)^{-k/2}\lesssim 1,$
since $k>2$. Using the above estimate, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{sm}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}=Q_{M}^{+}(|f_{l}|,|g_{k}|),$
Then, Theorem 3.1 implies
(4.17) $\|{\langle}v{\rangle}^{k}Q_{1}^{sm}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{l}\|_{L_{v}^{q_{1}}},$
since $p_{1}<2$.
Estimate for $Q_{1}^{r}$
Using the inequality $|u|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we bound
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|\geq|v_{1}^{*}|/10}\mathds{1}_{|v^{*}|>E/2}\lesssim(1+E)^{\frac{1-l}{2}}\lesssim 1,$
since $l>2$. Arguing as in the previous case, we obtain
(4.18) $\|{\langle}v{\rangle}^{k}Q_{1}^{r}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{1}}}\|g_{l}\|_{L_{v}^{q_{1}}},$
since $p_{1}<2$.
We now estimate the singular operators $Q_{1}^{s_{1}}$ and $Q_{1}^{s_{2}}$.
Estimate for $Q_{1}^{s_{1}}$
Using the facts $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$ and $0<\gamma\leq 1$, we have
(4.19) $\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim\frac{|u|^{\gamma}{\langle}v{\rangle}^{k-1}}{{\langle}v_{1}^{*}{\rangle}^{k}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}.$
We take different cases for $1<p_{2}<2$ and $p_{2}=1$.
$\bullet$ Case $1<p_{2}<2$ : Using (4.19) and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p_{2}}}\right)\left(\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p_{2}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.20) $\lesssim I_{1}^{1/q_{2}}(v)I_{2}^{1/p_{2}}(v),$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q_{2}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q_{2}/p_{2}}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{p_{2}}|g_{k}(v_{1}^{*})|^{lp_{2}}||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp_{2}}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Since $l>2$, (2.26) implies $I_{2}\lesssim{\langle}v{\rangle}^{p_{2}-2}\|g_{k}\|_{L_{v}^{p_{2}}}^{p_{2}}$.
In order to estimate $I_{1}$, we use the definition (4.15) of $\mathcal{B}_{1}(v)$, followed by (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ to write
$I_{1}(v)$ $\leq\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(u))|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(y))|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound, to obtain
$I_{1}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q_{2}/p_{2}}}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)}{\cos^{2}\theta(\sin\theta)^{\frac{2q_{2}}{p_{2}}-1}}\,d\theta\,d\nu$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)\cos\theta}{\cos^{3}\theta(\sin\theta)^{\frac{2q_{2}}{p_{2}}-1}}\,d\theta\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/4}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\cos\theta}{(\sin\theta)^{\frac{2q_{2}}{p_{2}}-1}}\,d\theta\,d\nu$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{{\langle}v{\rangle}^{-1/2}}^{1/\sqrt{2}}x^{1-\frac{2q_{2}}{p_{2}}}\,dx\,d\nu$
$\lesssim{\langle}v{\rangle}^{\frac{q_{2}}{p_{2}}-1}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\,d\nu={\langle}v{\rangle}^{\frac{q_{2}}{p_{2}}-1}\|f_{l}\|_{L_{v}^{q_{2}}}^{q_{2}},$
where we used the substitution $x:=\sin\theta$, and the fact that $p_{2}<2<q_{2}$ to estimate the integral in $x$. Note that the right-hand side of the above estimate is growing as $|v|\to\infty$, however the decay of $I_{2}$ is sufficient to counteract this growth.
Indeed, combining (4.20) with the estimates for $I_{1},I_{2}$ and using the exponents’ condition, we obtain
(4.21) ${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim|v|^{\frac{1}{p_{2}}-\frac{1}{q_{2}}+1-\frac{2}{p_{2}}}\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}}=\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}},\quad 1<p_{2}<2.$
$\bullet$ Case $p_{2}=1$ : We use again (4.19) and the definition of $\mathcal{B}_{1}(v)$ followed by (2.26) to bound
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\|f_{l}\|_{L_{v}^{\infty}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})|}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|\geq{\langle}v{\rangle}^{-1/2}}\,d\sigma\,dv_{1}$
$\leq\|f_{l}\|_{L_{v}^{\infty}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Since $l>2$, (2.26) implies
(4.22) ${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|\lesssim\|f_{l}\|_{L_{v}^{\infty}}\|g_{k}\|_{L_{v}^{1}}.$
Combining (4.21), (4.22), we conclude
(4.23) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}},\quad 1\leq p_{2}<2.$
Estimate for $Q_{1}^{s_{2}}$
Using (4.19) and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q_{1}}}\right)\left(\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q_{1}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<|v|^{-1/2}}\,d\sigma\,dv_{1}$
(4.24) $\lesssim I_{1}^{1/p_{1}}(v)I_{2}^{1/q_{1}}(v),$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{p_{1}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{1}/q_{1}}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{q_{1}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq_{1}}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Since $l>2$, (2.26) yields
$I_{2}\lesssim{\langle}v{\rangle}^{q_{1}-2}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}}.$
Notice again that the right-hand side of this estimate grows as $|v|\to\infty$ since $q_{1}>2$. However, we will see that the estimate for $I_{1}$ provides sufficient decay to counteract this growth.
Indeed, in order to estimate $I_{1}$, we use the same argument as for $Q_{1}^{s_{1}}$ using the definition (4.16) of $\mathcal{B}_{2}(v)$ instead (which imposes $|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<{\langle}v{\rangle}^{-1/2}$) and exchanging the exponents. We obtain
$I_{1}(v)\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{0}^{{\langle}v{\rangle}^{-1/2}}x^{1-\frac{2p_{1}}{q_{1}}}\,dx\,d\nu$ $\lesssim{\langle}v{\rangle}^{\frac{p_{1}}{q_{1}}-1}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\,d\nu$
$={\langle}v{\rangle}^{\frac{p_{1}}{q_{1}}-1}\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}},$
where we used the fact that $p_{1}<2<q_{1}$ for the convergence of the integral in $x$.
Combining (4.24) with the bounds for $I_{1},I_{2}$, and using the exponents’ condition, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)(v)|$ $\lesssim|v|^{\frac{1}{q_{1}}-\frac{1}{p_{1}}+1-\frac{2}{q_{1}}}\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}=\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}},$
therefore
(4.25) $\|Q_{1}^{s_{2}}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}.$
Conclusion of the proof
Combining (4.14) with (4.17), (4.18), (4.23), (4.25), estimate (4.13) follows.
4.3. The $L_{v}^{r}$– estimate for $Q_{0}$
Proposition 4.3.
Let $0<\gamma\leq 1$ and $2\leq r<\infty$. Consider $1\leq p_{0}\leq 2\leq q_{0}\leq r$ with $\frac{1}{p_{0}}+\frac{1}{q_{0}}=1+\frac{1}{r}$. Then, for any $k\geq l>2$ there holds the estimate
(4.26) $\|{\langle}v{\rangle}^{k}Q_{0}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p_{0}}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q_{0}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{k}:={\langle}v{\rangle}^{k}f$, $g_{l}:={\langle}v{\rangle}^{l}g$. We note that, since $p_{0}\leq 2\leq q_{0}$, we have $q_{0}^{\prime}\leq 2\leq p_{0}^{\prime}$ where $p_{0}^{\prime},q_{0}^{\prime}$ are the conjugate exponents of $p_{0},q_{0}$ respectively.
We use the decomposition (4.5) from Proposition 4.1.
Estimate for $Q_{0}^{sm},Q_{0}^{r}$
The operators $Q_{0}^{sm}$, $Q_{0}^{r}$ can be treated similarly to Proposition 4.1, we just use Theorem 3.2 instead of Theorem 3.1. Omitting details, we obtain
(4.27) $\|{\langle}v{\rangle}^{k}Q_{0}^{sm}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{0}}}\|g_{k}\|_{L_{v}^{q_{0}}},$
(4.28) $\|{\langle}v{\rangle}^{k}Q_{0}^{r}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{0}}}\|g_{k}\|_{L_{v}^{q_{0}}}.$
Estimate for $Q_{0}^{s}$
We take different cases for $1<p_{0}\leq 2\leq q_{0}<r$ and $p_{0}=1,\,q_{0}=r$.
$\bullet$ Case $1<p_{0}\leq 2\leq q_{0}<r$ : Using (4.8) and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})g_{l}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}|f_{k}(v^{*})|^{p_{0}/q_{0}^{\prime}}}{{\langle}v_{1}^{*}{\rangle}^{l}}\right)\left(|g_{l}(v_{1}^{*})|^{q_{0}/p_{0}^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}\right)\left(|f_{k}(v^{*})|^{p_{0}/r}|g_{l}(v_{1}^{*})|^{q_{0}/r}\right)$
$\hskip 142.26378pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.29) $\leq I_{1}^{1/q_{0}^{\prime}}(v)I_{2}^{1/p_{0}^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})|^{p_{0}}|g_{l}(v_{1}^{*})|^{q_{0}}\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{q_{0}^{\prime}}}{{\langle}v_{1}^{*}{\rangle}^{lq_{0}^{\prime}}}|f_{k}(v^{*})|^{p_{0}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g_{l}(v_{1}^{*})|^{q_{0}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{0}^{\prime}}\,d\sigma\,dv_{1}.$
By (2.25) we have $I_{1}(v)\leq{\langle}v{\rangle}^{q_{0}^{\prime}-2}\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}\leq\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}$, since $q_{0}^{\prime}\leq 2$.
For $I_{2}$, we use (2.7) and the substitution $y:=u=v-v_{1}$ to write
$I_{2}(v)$ $=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|g_{l}(v-R_{\sigma}^{-}(u))|^{q_{0}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{0}^{\prime}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|g_{l}(v-R_{\sigma}^{-}(y))|^{q_{0}}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|^{2p_{0}^{\prime}}\,dy\,d\sigma.$
We now use Proposition 2.1 to substitute $\nu=R_{\sigma}^{-}(y)$ and obtain
$I_{2}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|g_{l}(v-\nu)|^{q_{0}}|\widehat{\nu}\cdot\sigma|^{2(p_{0}^{\prime}-1)}\,d\nu\,d\sigma\lesssim\int_{{\mathbb{R}}^{3}}|g_{l}(v-\nu)|^{q_{0}}\,d\nu=\|g_{l}\|_{L_{v}^{q_{0}}}^{q_{0}}.$
Combining (4.29) with the estimates for $I_{1},I_{2}$, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|\lesssim\|f_{k}\|_{L^{p_{0}}_{v}}^{p_{0}/q_{0}^{\prime}}\|g_{l}\|_{L^{q_{0}}_{v}}^{q_{0}/p_{0}^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})|^{p_{0}}|g_{l}(v_{1}^{*})|^{q_{0}}\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above bound to the $r$-th power, integrating in $v$, and using (2.28), we obtain
$\|$ ${\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L^{r}_{v}}^{r}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}^{r-p_{0}}\|g_{l}\|_{L_{v}^{q_{0}}}^{r-q_{0}}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})|^{p_{0}}|g_{l}(v_{1}^{*})|^{q_{0}}\,d\sigma\,dv_{1}\,dv$
(4.30) $=\|f_{k}\|_{L_{v}^{p_{0}}}^{r-p_{0}}\|g_{l}\|_{L_{v}^{q_{0}}}^{r-q_{0}}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v)|^{p_{0}}|g_{l}(v_{1})|^{q_{0}}\,d\sigma\,dv_{1}\,dv\approx\|f_{k}\|_{L^{p_{0}}_{v}}^{r}\|g_{l}\|_{L^{q_{0}}_{v}}^{r}.$
$\bullet$ Case $p_{0}=1,\,q_{0}=r$ : We use again (4.8) and Hölder’s inequality, to bound instead as follows
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})g_{l}(v_{1}^{*})\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})|^{1/r^{\prime}}\right)\left(|f_{k}(v^{*})|^{1/r}|g_{l}(v_{1}^{*})|\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.31) $\lesssim\widetilde{I}_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})||g_{l}(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\right)^{1/r},$
where
$\widetilde{I}_{1}(v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{r^{\prime}}}{{\langle}v_{1}^{*}{\rangle}^{lr^{\prime}}}|f_{k}(v^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Since $l>2$ and $r\geq 2$, (2.25) implies $\widetilde{I}_{1}(v)\lesssim{\langle}v{\rangle}^{r^{\prime}-2}\|f_{k}\|_{L_{v}^{1}}\leq\|f_{k}\|_{L_{v}^{1}}$. Combining (4.31) with the bound on $\widetilde{I}_{1}$, we obtain
(4.32) ${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|\lesssim\|f_{k}\|_{L_{v}^{1}}^{1/r^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})||g_{l}(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above bound to the $r$-th power, integrating in $v$, and using (2.28), we obtain
$\|$ ${\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L^{r}_{v}}^{r}\lesssim\|f_{k}\|_{L^{1}_{v}}^{r-1}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})||g_{l}(v_{1}^{*})|\,d\sigma\,dv_{1}\,dv$
(4.33) $=\|f_{k}\|_{L^{1}_{v}}^{r-1}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v)||g_{l}(v_{1})|\,d\sigma\,dv_{1}\,dv\approx\|f_{k}\|_{L^{1}_{v}}^{r}\|g_{l}\|_{L^{r}_{v}}^{r}.$
Combining (4.30), (4.33), we deduce
(4.34) $\|{\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L^{r}_{v}}\lesssim\|f_{k}\|_{L^{p_{0}}_{v}}\|g_{l}\|_{L^{q_{0}}_{v}},\quad 1\leq p_{0}\leq 2\leq q_{0}\leq r.$
Conclusion of the proof
Combining (4.5) with (4.27), (4.28), (4.34), estimate (4.26) follows.
4.4. The $L_{v}^{r}$– estimate for $Q_{1}$
Proposition 4.4.
Let $0<\gamma\leq 1$ and $2<r<\infty$. Consider $1<p_{1}<2<q_{1}<r$ and $1\leq p_{2}<2<q_{2}\leq r$ with $\frac{1}{p_{1}}+\frac{1}{q_{1}}=\frac{1}{p_{2}}+\frac{1}{q_{2}}=1+\frac{1}{r}$. Then for any $k\geq l>2$, there holds the estimate
(4.35) $\|{\langle}v{\rangle}^{k}Q_{1}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p_{1}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q_{1}}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q_{2}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p_{2}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$. We use the decomposition (4.14) from Proposition 4.2.
Estimate for $Q_{1}^{sm}$, $Q_{1}^{r}$
The operators $Q_{1}^{sm},Q_{1}^{r}$ can be treated similarly to $Q_{1}^{sm},Q_{1}^{r}$ from Proposition 4.2, we just use Theorem 3.2 instead of Theorem 3.1. Omitting details, we obtain
(4.36) $\|{\langle}v{\rangle}^{k}Q_{1}^{sm}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}},$
(4.37) $\|{\langle}v{\rangle}^{k}Q_{1}^{r}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}.$
We now estimate the singular operators $Q_{1}^{s_{1}}$ and $Q_{1}^{s_{2}}$.
Estimate for $Q_{1}^{s_{1}}$
We will take different cases for $1<p_{2}<2<q_{2}<r$ and $p_{2}=1,\,q_{2}=r$.
$\bullet$ Case $1<p_{2}<2<q_{2}<r$ : Since $q_{2}>2$, we have $q_{2}^{\prime}<2$, thus
(4.38) $\frac{1}{r}$ $=1-\frac{1}{p_{2}^{\prime}}-\frac{1}{q_{2}^{\prime}}=1-\frac{2}{q_{2}^{\prime}}+\frac{1}{q_{2}^{\prime}}-\frac{1}{p_{2}^{\prime}}<\frac{1}{q_{2}^{\prime}}-\frac{1}{p_{2}^{\prime}},$
Pick $\alpha>1$ with $\frac{\alpha}{r}<\frac{1}{q^{\prime}_{2}}-\frac{1}{p^{\prime}_{1}}$, which is possible to pick due to (4.38). Using (4.19) and Hölder’s inequality, we bound
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|^{q_{2}/p_{2}^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\right)\left(\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q_{2}^{\prime}}|g_{k}(v_{1}^{*})|^{p_{2}/q_{2}^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left(\frac{|f_{l}(v^{*})|^{q_{2}/r}|g_{k}(v_{1}^{*})|^{p_{2}/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/r}}\right)\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.39) $\leq I_{1}^{1/p_{2}^{\prime}}(v)I_{2}^{1/q_{2}^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q_{2}}|g_{k}(v_{1}^{*})|^{p_{2}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q_{2}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{q_{2}^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}|g_{k}(v_{1}^{*})|^{p_{2}}}{{\langle}v^{*}{\rangle}^{lq^{\prime}_{2}}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
By (2.26) we have $I_{2}(v)\lesssim{\langle}v{\rangle}^{q_{2}^{\prime}-2}\|g_{k}\|_{L_{v}^{p_{2}}}^{p_{2}}$.
For $I_{1}$, we use the definition (4.15) of $\mathcal{B}_{1}(v)$, followed by (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$, to write
$I_{1}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f_{l}(v-R_{\sigma}^{+}(u))|^{q_{2}}\mathds{1}_{|v|>10}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\,dv_{1}\,d\sigma,$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f_{l}(v-R_{\sigma}^{+}(y))|^{q_{2}}\mathds{1}_{|v|>10}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound and obtain
$I_{1}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)}{\cos^{2}\theta\,(\sin\theta)^{2p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})-1}}\,d\theta\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)\cos\theta}{\cos^{3}\theta\,(\sin\theta)^{2p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})-1}}\,d\theta\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{0}^{\pi/4}\frac{\mathds{1}_{\sin\theta\geq{\langle}v{\rangle}^{-1/2}}\cos\theta}{(\sin\theta)^{2p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})-1}}\,d\theta\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\mathds{1}_{|v|>10}\int_{{\langle}v{\rangle}^{-1/2}}^{1/\sqrt{2}}x^{1-2p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}\,dx\,d\nu$
$\lesssim{\langle}v{\rangle}^{p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})-1}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q_{2}}\,d\nu={\langle}v{\rangle}^{p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})-1}\|f_{l}\|_{L_{v}^{q_{2}}}^{q_{2}},$
where we used the substitution $x:=\sin\theta$, and the fact that $\frac{\alpha}{r}<\frac{1}{q_{2}^{\prime}}-\frac{1}{p_{2}^{\prime}}$ to estimate the integral in $x$.
Combining (4.39) with the bounds on $I_{1},I_{2}$ and using the exponents’ condition, we obtain the point-wise bound
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim{\langle}v{\rangle}^{\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r}-\frac{1}{p_{2}^{\prime}}+1-\frac{2}{q_{2}^{\prime}}}\|f_{l}\|_{L_{v}^{q_{2}}}^{q_{2}/p_{2}^{\prime}}\|g_{k}\|_{L_{v}^{p_{2}}}^{p_{2}/q_{2}^{\prime}}$
$\hskip 28.45274pt\times\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r}$
$\hskip-56.9055pt={\langle}v{\rangle}^{\frac{1-\alpha}{r}}\|f_{l}\|_{L_{v}^{q_{2}}}^{q_{2}/p_{2}^{\prime}}\|g_{k}\|_{L_{v}^{p_{2}}}^{p_{2}/q_{2}^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above bound to the $r$-th power and integrating in $v$, we obtain
$\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{r}}^{r}$ $\lesssim\|f_{l}\|_{L_{v}^{q_{2}}}^{r-q_{2}}\|g_{k}\|_{L_{v}^{p_{2}}}^{r-p_{2}}$
(4.40) $\hskip 14.22636pt\times\left(\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v{\rangle}^{1-\alpha}\frac{|f_{l}(v^{*})|^{q_{2}}|g_{k}(v_{1}^{*})|^{p_{2}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv\right).$
Now, we observe that
(4.41) $\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|<|v_{1}^{*}|/10}\chi(\widehat{u}\cdot\sigma)\leq\mathds{1}_{E>100}\mathds{1}_{|v|^{2}>\frac{E^{1/2}}{6}}\chi(\widehat{u}\cdot\sigma).$
Indeed, consider $|v|>10$ and $(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}$ with $|v_{1}^{*}|^{2}\geq E/2$, $|v^{*}|<|v_{1}^{*}|/10$ and $\widehat{u}\cdot\sigma>0$. By (2.11), (2.12), (1.6) and the triangle inequality, we have
$|R_{\sigma}^{-}(u)|=|u||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|=|u|\sqrt{\frac{1-\widehat{u}\cdot\sigma}{2}}<\frac{|u|}{\sqrt{2}}\leq\frac{|v_{1}^{*}|+|v^{*}|}{\sqrt{2}}\leq\frac{11|v_{1}^{*}|}{10\sqrt{2}}.$
Then, (2.7) followed by the triangle inequality and the above bound implies
$|v|=|v_{1}^{*}+R_{\sigma}^{-}(u)|\geq|v_{1}^{*}|-|R_{\sigma}^{-}(u)|>\left(1-\frac{11}{10\sqrt{2}}\right)|v_{1}^{*}|>\frac{|v_{1}^{*}|}{5}.$
We also have $E=|v^{*}|^{2}+|v_{1}^{*}|^{2}<\frac{11}{10}|v_{1}^{*}|^{2}$, which combined with the above bound yields $|v|>\frac{E^{1/2}}{6}$. Since $|v|>10$, we also have $100<|v|^{2}\leq E$, and (4.41) follows.
Recalling (4.15), and using (4.41) we obtain
(4.42) $\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\leq\mathds{1}_{E>100}\mathds{1}_{|v|>\frac{E^{1/2}}{6}}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}>6E^{-1/2}}\chi(\widehat{u}\cdot\sigma).$
Using (4.42), followed by (2.12) and (2.28), the integral term of (4.40) can be estimated as follows
$\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}$ ${\langle}v{\rangle}^{1-\alpha}\frac{|f_{l}(v^{*})|^{q_{2}}|g_{k}(v_{1}^{*})|^{p_{2}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$
$\leq\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v{\rangle}^{1-\alpha}\frac{|f_{l}(v^{*})|^{q_{2}}|g_{k}(v_{1}^{*})|^{p_{2}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{E>100}\mathds{1}_{|v|>\frac{E^{1/2}}{6}}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}>6E^{-1/2}}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$
$\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{\frac{1-\alpha}{2}}\frac{|f_{l}(v^{*})|^{q_{2}}|g_{k}(v_{1}^{*})|^{p_{2}}}{\left(1-\widehat{u}\cdot\sigma\right)^{\alpha}}\mathds{1}_{E>100}\mathds{1}_{0<\widehat{u}\cdot\sigma<1-12E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q_{2}}|g_{k}(v_{1})|^{p_{2}}\mathds{1}_{E>100}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{0<\widehat{u}\cdot\sigma\leq 1-12E^{-1/2}}}{\left(1-\widehat{u}\cdot\sigma\right)^{\alpha}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q_{2}}|g_{k}(v_{1})|^{p_{2}}\mathds{1}_{E>100}\int_{0}^{1-12E^{-1/2}}\frac{1}{\left(1-x\right)^{\alpha}}\,dx\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q_{2}}|g_{k}(v_{1})|^{p_{2}}\mathds{1}_{E>100}\int_{12E^{-1/2}}^{1}z^{-\alpha}\,dz\,dv_{1}\,dv$
(4.43) $\lesssim\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{q_{2}}|g_{k}(v_{1})|^{p_{2}}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{q_{2}}}^{q_{2}}\|g_{k}\|_{L_{v}^{p_{2}}}^{p_{2}},$
where we used the fact that $\alpha>1$ to estimate the integral in $z$.
Combining (4.40), (4.43), we obtain
(4.44) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}},\quad 1<p_{2}<2<q_{2}<r.$
$\bullet$ Case $p_{2}=1,\,q_{2}=r$ : We use again (4.19) and Hölder’s inequality to bound instead as follows
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})|^{1/r^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/r^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)\left(\frac{|f_{l}(v^{*})||g_{k}(v_{1}^{*})|^{1/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/r^{\prime}}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.45) $\leq\widetilde{I}_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{r}|g_{k}(v_{1}^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2(r-1)}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r},$
where
$\widetilde{I}_{1}(v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{r^{\prime}}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lr^{\prime}}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
By (2.26), we have $\widetilde{I}_{1}(v)\lesssim{\langle}v{\rangle}^{r^{\prime}-2}\|g_{k}\|_{L_{v}^{1}}$. Combining the bound on $\widetilde{I}_{1}$ with (4.31), raising to the $r$-th power, and integrating in $v$, we obtain
(4.46) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{r}}^{r}\lesssim\|g_{k}\|_{L_{v}^{1}}^{r-1}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v{\rangle}^{1-\alpha}\frac{|f_{l}(v^{*})|^{r}|g_{k}(v_{1}^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv,$
where we write $\alpha:=\frac{r}{r^{\prime}}=r-1>1$, since $r>2$.
Now we argue identically as in the previous case to obtain
(4.47) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{r}}\|g_{k}\|_{L_{v}^{1}}.$
Combining (4.44), (4.47), we deduce
(4.48) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{1}}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{q_{2}}}\|g_{k}\|_{L_{v}^{p_{2}}},\quad 1\leq p_{2}<2<q_{2}\leq r.$
Estimate for $Q_{1}^{s_{2}}$
Using the facts $|u|\leq E^{1/2}$, $0<\gamma\leq 1$, and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)|$ $\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|^{p_{1}/q_{1}^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p_{1}^{\prime}}}\right)\left(\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p_{1}^{\prime}}|g_{k}(v_{1}^{*})|^{q_{1}/p_{1}^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left(|f_{l}(v^{*})|^{p_{1}/r}|g_{k}(v_{1}^{*})|^{q_{1}/r}\right)\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.49) $\leq I_{1}^{1/q_{1}^{\prime}}(v)I_{2}^{1/p_{1}^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{p_{1}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q_{1}^{\prime}/p_{1}^{\prime}}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{p_{1}^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}|g_{k}(v_{1}^{*})|^{q_{1}}}{{\langle}v^{*}{\rangle}^{lp_{1}^{\prime}}}\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
By (2.26), we have $I_{2}(v)\lesssim{\langle}v{\rangle}^{p_{1}^{\prime}-2}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}}$.
For $I_{1}$, we use the definition (4.16) of $\mathcal{B}_{2}$, followed (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ to write
$I_{1}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f_{l}(v-R_{\sigma}^{+}(u))|^{p_{1}}\mathds{1}_{|v|>10}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q_{1}^{\prime}/p_{1}^{\prime}}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f_{l}(v-R_{\sigma}^{+}(y))|^{p_{1}}\mathds{1}_{|v|>10}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}>{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q_{1}^{\prime}/p_{1}^{\prime}}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound and obtain
$I_{1}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q_{1}^{\prime}/p_{1}^{\prime}}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q_{1}^{\prime}/p_{1}^{\prime}}}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta<{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)}{\cos^{2}\theta\,(\sin\theta)^{\frac{2q_{1}^{\prime}}{p_{1}^{\prime}}-1}}\,d\theta\,d\nu$
$=\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{0}^{\pi/2}\frac{\mathds{1}_{\sin\theta<{\langle}v{\rangle}^{-1/2}}\chi(2\cos^{2}\theta-1)\cos\theta}{\cos^{3}\theta\,(\sin\theta)^{\frac{2q_{1}^{\prime}}{p_{1}^{\prime}}-1}}\,d\theta\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{0}^{\pi/4}\frac{\mathds{1}_{\sin\theta<{\langle}v{\rangle}^{-1/2}}\cos\theta}{(\sin\theta)^{\frac{2q_{1}^{\prime}}{p_{1}^{\prime}}-1}}\,d\theta\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\mathds{1}_{|v|>10}\int_{0}^{{\langle}v{\rangle}^{-1/2}}x^{1-\frac{2q_{1}^{\prime}}{p_{1}^{\prime}}}\,dx\,d\nu$
$\lesssim{\langle}v{\rangle}^{\frac{q_{1}^{\prime}}{p_{1}^{\prime}}-1}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p_{1}}\,d\nu={\langle}v{\rangle}^{\frac{q_{1}^{\prime}}{p_{1}^{\prime}}-1}\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}},$
where we used the substitution $x:=\sin\theta$ and the fact that $q^{\prime}_{1}<p^{\prime}_{1}$ for the convergence of the integral in $x$.
Combining the bounds on $I_{1},I_{2}$ with (4.49), for $1<p_{1}<2<q_{1}<r$, we have
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)|$ $\lesssim{\langle}v{\rangle}^{1/r}\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}/q_{1}^{\prime}}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}/p^{\prime}_{1}}$
(4.50) $\quad\times\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above inequality to the $r$-th power and integrating in $v$, we obtain
$\|{\langle}v{\rangle}^{k}Q_{1}^{s_{2}}(f,g)\|_{L_{v}^{r}}^{r}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}^{r-p_{1}}\|g_{k}\|_{L_{v}^{q_{1}}}^{r-q_{1}}$
(4.51) $\quad\times\left(\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v{\rangle}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv\right)$
Recalling (4.16), and using (4.41) we obtain
(4.52) $\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\leq\mathds{1}_{E>100}\mathds{1}_{|v|>\frac{E^{1/2}}{6}}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}<6E^{-1/2}}\chi(\widehat{u}\cdot\sigma).$
Using (4.52), followed by (2.12) and (2.28), the integral term of (4.51) can be estimated as
$\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}$ ${\langle}v{\rangle}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$
$\leq\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{E>100}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}<6E^{-1/2}}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{E>100}\mathds{1}_{\widehat{u}\cdot\sigma>1-12E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p_{1}}|g_{k}(v_{1})|^{q_{1}}\mathds{1}_{E>100}\int_{{\mathbb{S}}^{2}}\mathds{1}_{\widehat{u}\cdot\sigma>1-12E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p_{1}}|g_{k}(v_{1})|^{q_{1}}\mathds{1}_{E>100}\int_{1-12E^{-1/2}}^{1}\,dx\,dv_{1}\,dv$
(4.53) $\approx\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{p_{1}}|g_{k}(v_{1})|^{q_{1}}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}}.$
By (4.51), (4.53), we conclude
(4.54) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{2}}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}.$
Conclusion of the proof
Combining (4.14) with (4.36), (4.48), (4.54), estimate (4.35) follows.
4.5. Proof of Theorem 1.1
We conclude with the proof of Theorem 1.1. Using decomposition (4.1), the triangle inequality and Propositions 4.14.4, we obtain
(4.55) $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},$
for all $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Recalling the definition of $\|\cdot\|_{\bm{\mathcal{X}_{p,q}^{k}}}$ from (1.19), the above bound implies
(4.56) $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}\|g\|_{\bm{\mathcal{X}_{p,q}^{k}}},\quad\forall f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
By Lemma A.1, we can uniquely extend $Q^{+}$ by density to a bounded bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k}L^{r}_{v}$ by a standard density argument.
Finally, we prove that (1.21) holds for any $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Indeed, consider $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Using Lemma A.1, we pick sequences $\{f_{n}\}_{n=1}^{\infty}$, $\{g_{n}\}_{n=1}^{\infty}\subset C_{c}^{\infty}({\mathbb{R}}^{3})$ with $f_{n}\to f$ and $g_{n}\to g$ as $n\to\infty$ in $\bm{\mathcal{X}_{p,q}^{k}}$. Then, by (4.55), for any $n\in\mathbb{N}$ we have
$\|{\langle}v{\rangle}^{k}Q^{+}(f_{n},g_{n})\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f_{n}\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{l}g_{n}\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{p}}.$
Letting $n\to\infty$ in the above bound, and using continuity of $Q^{+}$ for the left-had side as well as (1.20) for the right-hand side, estimate (1.21) follows. The proof is complete.
Appendix A
In this appendix, we record a basic density result.
Lemma A.1.
Let $k\geq 0$ and $1\leq p,q<\infty$. Then $C_{c}^{\infty}({\mathbb{R}}^{3})$ is a dense subspace of $\bm{\mathcal{X}_{p,q}^{k}}$.
Proof.
Let $f\in\bm{\mathcal{X}_{p,q}^{k}}$ and $\epsilon>0$. We will construct $\psi\in C_{c}^{\infty}({\mathbb{R}}^{3})$ with $\|\psi-f\|_{\bm{\mathcal{X}_{p,q}^{k}}}<\epsilon$.
Define $\widetilde{f}={\langle}v{\rangle}^{k}f\in L^{p}_{v}\cap L^{q}_{v}$, and given $R>0$ let us write $\widetilde{f}_{R}:=\widetilde{f}\,\mathds{1}_{|x|\leq R}$. Since $\widetilde{f}\in L^{p}_{v}\cap L^{q}_{v}$, the dominated convergence theorem implies that $\widetilde{f}_{R}\to\widetilde{f}$ as $R\to\infty$ both in $L^{p}_{v}$ and $L^{q}_{v}$, thus there exists $R_{0}>0$ such that
(A.1) $\|\widetilde{f}_{R_{0}}-\widetilde{f}\|_{L^{p}_{v}}<\frac{\epsilon}{2},\quad\|\widetilde{f}_{R_{0}}-\widetilde{f}\|_{L^{q}_{v}}<\frac{\epsilon}{2}.$
Now consider a mollifier $(\rho_{\delta})_{0<\delta\leq 1}$. Then, $\rho_{\delta}\ast\widetilde{f}_{R_{0}}\in C_{c}^{\infty}({\mathbb{R}}^{3})$ and $\rho_{\delta}\ast\widetilde{f}_{R_{0}}\to\widetilde{f}_{R_{0}}$ as $\delta\to 0$ both in $L^{p}_{v}$ and $L^{q}_{v}$. Thus there exists $0<\delta_{0}<1$ such that
(A.2) $\|\rho_{\delta_{0}}\ast\widetilde{f}_{R_{0}}-\widetilde{f}_{R_{0}}\|_{L^{p}_{v}}<\frac{\epsilon}{2},\quad\|\rho_{\delta_{0}}\ast\widetilde{f}_{R_{0}}-\widetilde{f}_{R_{0}}\|_{L^{q}_{v}}<\frac{\epsilon}{2}.$
Define $\widetilde{\psi}:=(\rho_{\delta_{0}}\ast\widetilde{f}_{R_{0}})\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Then by the triangle inequality and (A.1), (A.2), we obtain
$\|\widetilde{\psi}-\widetilde{f}\|_{L^{p}_{v}}<\epsilon,\quad\|\widetilde{\psi}-\widetilde{f}\|_{L^{q}_{v}}<\epsilon.$
Defining $\psi:={\langle}v{\rangle}^{-k}\widetilde{\psi}\in C_{c}^{\infty}({\mathbb{R}}^{3})$, we have
$\|\psi-f\|_{\bm{\mathcal{X}_{p,q}^{k}}}$ $=\max\left\{\|{\langle}v{\rangle}^{k}(\psi-f)\|_{L^{p}_{v}},\|{\langle}v{\rangle}^{k}(\psi-f)\|_{L^{q}_{v}}\right\}$
$=\max\left\{\|\widetilde{\psi}-\widetilde{f}\|_{L^{p}_{v}},\|\widetilde{\psi}-\widetilde{f}\|_{L^{q}_{v}}\right\}<\epsilon,$
and the claim follows.