### v4

# Convolution estimates for the Boltzmann gain operator with hard-spheres

Abstract.
We prove new moment-preserving polynomially weighted convolution estimates for the gain operator of the Boltzmann equation with hard potentials, including the critical case of hard-spheres. Our approach relies crucially on a novel cancellation mechanism dealing with the pathological case of energy-absorbing collisions (that is, collisions that accumulate energy to only one of the outgoing particles). This difficulty is specific to hard potentials, and is not present for Maxwell molecules. Our method quantifies the heuristic that, while energy-absorbing collisions occur with non-trivial probability, they are statistically rare, and therefore do not affect the overall averaging behavior of the gain operator. At the technical level, our proof relies solely on tools from kinetic theory, such as geometric identities and angular averaging.
1. Introduction
1.1. Problem setup and goal of this paper
The Boltzmann equation is the central equation of collisional kinetic theory and one of the most celebrated equations of mathematical physics. It describes dilute gases of microscopic interacting particles at the mesoscopic/kinetic level; more precisely, it predicts the time evolution of the probability density $f(t,x,v)$ of a dilute gas in non-equilibrium. The Boltzmann equation reads
(1.1) $\partial_{t}f+v\cdot\nabla_{x}f=Q(f,f),$
where the advection operator $\partial_{t}+v\cdot\nabla_{x}$ accounts for the rectilinear motion of particles, while $Q(f,f)$ is the quadratic collisional operator encoding the binary interactions between gas particles. It is given by
(1.2) $Q(f,f)(t,x,v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}B(u,\sigma)\Big(f(t,x,v^{*})f(t,x,v_{1}^{*})-f(t,x,v)f(t,x,v_{1})\Big)\,d\sigma\,dv_{1},$
where $v,v_{1}$ denote the pre-collisional velocities of the incoming particles, $v^{*}$, $v_{1}^{*}$ denote the post-collisional velocities, and $u:=v-v_{1}$ denotes the relative velocity of the particles before the collision. Simplifying the notation slightly, we drop the dependence of the arguments on $t,x.$
In the case of perfectly elastic collisions, and assuming without loss of generality that all particles are of equal mass $m=1$, the post-collisional velocities are given by:
(1.3) $\begin{split}v^{*}&=V-\frac{|u|}{2}\sigma,\quad v_{1}^{*}=V+\frac{|u|}{2}\sigma,\end{split}$
where $V:=\frac{v+v_{1}}{2}$ is the center of mass of the velocities of the incoming velocities and $\sigma\in{\mathbb{S}}^{2}$ represents the scattering direction of the collision. It is then straightforward to see that the momentum and energy are conserved, as well as the relative velocity magnitude:
(1.4) $v^{*}+v_{1}^{*}$ $=v+v_{1},$
(1.5) $|v^{*}|^{2}+|v_{1}^{*}|^{2}$ $=|v|^{2}+|v_{1}|^{2},$
(1.6) $|v^{*}-v_{1}^{*}|$ $=|v-v_{1}|.$
The kernel $B(u,\sigma)$ appearing in (1.2) is the collisional cross-section, which is assumed to be of the form
(1.7) $B(u,\sigma)=|u|^{\gamma}\,b(\widehat{u}\cdot\sigma),\quad-3<\gamma\leq 1,$
where the angular cross-section $b:[-1,1]\to{\mathbb{R}}$ satisfies:
(H.1) $b(z)$ $\geq 0,\quad\forall\,z\in[-1,1],$
(H.2) $b(-z)$ $=b(z),\quad\forall\,z\in[-1,1],$
(H.3) $b\in L^{\infty}([-1,1]).$
The parameter $\gamma$ in the radial part of the cross-section is of crucial importance: it determines the type of intramolecular potential considered.
  • The case $0<\gamma\leq 1$ corresponds to hard potentials, where the interaction is strong and of short range. They include the celebrated hard-sphere model ($\gamma=1,\,b=1$), where particles interact like impenetrable billiard balls. Hard-spheres are of particular importance since it was the model originally introduced by Boltzmann [8] and Maxwell [30]. It remains the only model for which the Boltzmann has been rigorously derived from deterministic many particle dynamics [27, 22, 15].
  • The case $\gamma=0$ corresponds to Maxwell molecules.
  • The case $3<\gamma<0$ corresponds to soft potentials. This type of interaction is smoother and of longer range.
Both physically and mathematically, these three types of potentials lead to different kinds of difficulties that typically require separate treatment. In this paper we focus on hard potentials, i.e. we assume $0<\gamma\leq 1$. In the course of the proof, we also improve existing results on Maxwell molecules with bounded cross-sections.
Regarding the assumptions on the angular cross-section $b$, condition (H.2) corresponds to indistinguishably of the particles, while the boundedness condition (H.3) is a slightly stronger assumption than the usual Grad’s cut-off, which is the mere integrability condition $\int_{{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)\,d\sigma<\infty.$ Yet (H.3) covers most cases of physical interest for hard potentials, and in particular the classical hard-sphere model ($\gamma=1,\,b=1$). Most importantly, since (H.3) is a special case of Grad’s cut-off, it implies that the collisional operator can be written in gain-loss form
(1.8) $Q(f,f)=Q^{+}(f,f)-Q^{-}(f,f),$
where the gain and loss operators are respectively given by
(1.9) $Q^{+}(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\,d\sigma\,dv_{1},$
(1.10) $Q^{-}(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)f(v_{1})\,d\sigma\,dv_{1}.$
Half-sphere restriction and bilinear operators
Exploiting spherical symmetry, it is often convenient to restrict the collisional operators to a half-sphere. For hard potentials, the most natural restriction is to fix the outgoing particle with the largest velocity (for Maxwell molecules we will restrict to the half-sphere corresponding to $\widehat{u}\cdot\sigma>0$ instead, see Section 3). Namely, by conservation of energy we have $|v^{*}|^{2}+|v_{1}^{*}|^{2}=|v|^{2}+|v_{1}|^{2}:=E$, thus either $|v_{1}^{*}|^{2}\geq E/2$ or $|v^{*}|^{2}\geq E/2$. Since the set where $|v_{1}^{*}|^{2}=|v^{*}|^{2}=E/2$ is of measure zero, we obtain
$Q^{+}(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(-\widehat{u}\cdot\sigma)f(v_{1}^{*})f(v^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1},$
where in the second part of the summation we used the substitution $\sigma\to-\sigma$ (which just interchanges $v^{*}$ and $v_{1}^{*}$ by (1.3)) and for the last equality we used (H.2). Similarly, we obtain
(1.11) $Q^{-}(f,f)=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)g(v_{1})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}.$
Motivated by this computation, we introduce the bilinear forms:
(1.12) $Q(f,g)$ $:=Q^{+}(f,g)-Q^{-}(f,g),$
(1.13) $Q^{+}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1},$
(1.14) $Q^{-}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)g(v_{1})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}.$
Clearly, the loss operator $Q^{-}(f,g)$ is much simpler to analyze, and essentially behaves like the product $f(|v|^{\gamma}\ast g).$
In contrast, the gain operator $Q^{+}(f,g)$ has a much richer structure due to the averaging over the angular variable and exhibits regularization effects [25, 28]. As a result, it is the main object to understand when developing the mathematical theory of the Boltzmann and related equations.
Quantifying these regularizing effects is subtle, though, because of the complex structure of the operator. The first step is to choose a suitable topology to measure the gain of regularity. The most relevant functional analytic setting is polynomially weighted Lebesgue spaces only in the velocity variable ${\langle}v{\rangle}^{-k}L^{p}_{v},$ for several reasons:
  • Such bounds are space translation invariant, thus they can be applied whether the spatial domain is bounded or unbounded. Indeed, ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$ estimates easily follow from Hölder’s inequality. This feature is important to address questions related to the hydrodynamic limit of (1.1). In that field, most mathematical results are proved in spatially bounded domains such as the torus. Without being exhaustive, we refer to the recent work of Deng-Hani-Ma [16] for a complete derivation of fundamental fluid dynamics equations from many-body classical dynamics. This followed their groundbreaking work [15] on the long-time derivation of the Boltzmann equation from hard-sphere dynamics which extended Lanford’s theorem [27, 22] globally in time.
  • Polynomial weights are the most suitable for studying the long-time behavior of the equation for large data, which is a major open problem, as outlined in [37]. Large data results for the space homogeneous (i.e. space independent) Boltzmann equation for hard potentials have been proved by establishing generation and propagation of moments [34, 20, 17, 38, 40, 11, 31], which correspond exactly to Lebesgue spaces with polynomial weights in $v$. Moreover, as noted by L. Desvillettes and C. Villani [18], the evolution cannot be treated by linearization techniques which typically require exponential weights.
  • Working with polynomial weights is also more natural in the closely related field of wave turbulence, where equilibria are typically inverse polynomials in $v$ (e.g. Rayleigh-Jeans distributions or the celebrated Kolmogorov-Zakharov spectra, see e.g. Nazarenko [33] for more details). Therefore, ideas from the present paper could also be helpful in studying wave kinetic equations, as shown in our previous work [3]. In an upcoming article [5] we are planning to use ideas and techniques of the current paper in order to prove optimal local well-posedness of the wave kinetic equation in $L^{p}$ spaces for any $2\leq p\leq\infty$. Such a result would extend the work of Germain-Ionescu-Tran [23] who obtained optimal well-posedness in $L^{2}$ and in $L^{\infty}$.
  • Finally, the conditions for the solution to be derivable from hard-sphere dynamics, and thus solve Hilbert’s sixth problem require weights only in $v$, see the statement of Theorem 1 in the seminal work of Deng-Hani-Ma [15].
In this paper, we use the averaging present in the gain operator to prove new moment-preserving convolution estimates in such polynomially weighted spaces. We work in the case of hard potentials, and the results include hard-spheres. Our approach relies crucially on a novel cancellation mechanism to deal with pathological collisions that accumulate energy to only one of the outgoing particles. As we shall see, this type of collisions distinguishes hard potentials from Maxwell molecules. Hence, a thorough understanding of these interactions is crucial for the study of hard potentials.
Related literature
After the foundational works of Carleman [13, 14] ($1<p<\infty$) and Arkeryd [6] ($p=\infty$) on the propagation of $L^{p}$–norms of solutions to the homogeneous Boltzmann equation, it was Gustafsson [24, 25] who first noticed the convolution structure of the gain collisional operator by proving a Young-type inequality for $Q^{+}(f,g)$, using an intricate nonlinear interpolation procedure. Later, in a pioneering work, Lions [28] used Fourier integral operators techniques to show that $Q^{+}(f,g)$ gives a derivative gain. Wennberg [39] simplified the proof of Lions and arrived at the same estimates by using the Carleman representation and classical Fourier analysis tools. See also the works [12, 29] for another family of estimates on the quadratic operator $Q^{+}(f,f)$ that are obtained by much simpler means. Based on these results, Toscani-Villani [36] and Mouhot-Villani [32] studied the propagation of $L^{p}$–norms and smoothness of solutions to the homogeneous Boltzmann equation. However, all these early results were obtained assuming point-wise truncations in the angular variable. More precisely, in these works singular collisions such as grazing $(\widehat{u}\cdot\sigma=0)$ or frontal $(\widehat{u}\cdot\sigma=\pm 1)$ are discarded.
More recently, Alonso-Carneiro [1] used Fourier analysis and radial symmetrization techniques to obtain a Young-type inequality for Maxwell molecules, under integrability assumptions on the cross-section $b$ that replaced point-wise truncations. The dependence on the cross-section is embedded in the constants of their estimates. Special cases of those estimates were previously obtained in [19] and [21]. Soon after that, Alonso-Carneiro-Gamba [2] extended the results of [1] to inelastic collisions, and also proved a Hardy-Littlewood-Sobolev type of inequality for soft potentials ($3<\gamma<0$).
We now focus on the most relevant prior results for our work, namely weighted convolution estimates for the gain operator without additional angular cut-offs.
Simple estimates of this kind can be straightforwardly obtained from the Maxwell molecule case, at the price of increasing weights on the arguments (see [1, 2]). These bounds are moment-increasing. More precisely for $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$, the estimate provided by [1, 2] is:
$\|{\langle}v{\rangle}^{k}\,Q^{+}(f,g)\|_{L^{r}_{v}}\lesssim\|{\langle}v{\rangle}^{k+\gamma}\,f\|_{L^{p}_{v}}\|{\langle}v{\rangle}^{k+\gamma}\,g\|_{L^{q}_{v}},\quad{\langle}v{\rangle}:=\sqrt{1+|v|^{2}},$
which is not sufficient to prove local/global well-posedness of the Boltzmann equation in spaces of the form ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$, or propagation of moments. The main difficulty is to offset the linear growth of the potential. Following Lions [28], a now classical (but highly technical) approach is based on decay properties of oscillatory integrals. Indeed, denoting $\widehat{f}$ and $\widehat{g}$ the Fourier transforms of $f$ and $g,$ (1.13) can be written (taking $b=1$ for simplicity)
$\int_{\mathbb{R}^{9}}\widehat{f}(\eta)\widehat{g}(\xi)e^{-iV\cdot(\xi+\eta)}|u|^{\gamma}\int_{{\mathbb{S}}^{2}}e^{-i\frac{|u|}{2}(\xi-\eta)\cdot\sigma}\,d\sigma\,dv_{1}d\xi d\eta.$
Standard stationary phase considerations (see e.g. [35]) give the asymptotic behavior of the inner integral for large $|u|$:
$\int_{{\mathbb{S}}^{2}}e^{-i\frac{|u|}{2}(\xi-\eta)\cdot\sigma}\,d\sigma\approx\frac{e^{\pm i|u||\xi-\eta|}}{|u||\xi-\eta|}.$
The decay in $|u|$ compensates for the growth of the potential. The remaining challenge is to prove $L^{p}$–boundedness of the resulting Fourier integral operator. A very recent paper [26] adapts this method to the hard potential case to produce a moment-preserving estimate for the gain operator with hard potentials. However, in the case of hard-spheres, this work only gives a single estimate, namely $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L^{6}_{v}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L^{1}_{v}}\|{\langle}v{\rangle}^{k}g\|_{L^{2}_{v}}$.
In comparison, we propose a kinetic alternative to the Fourier analytic approach, which we present now. It allows us to prove a much wider set of estimates.
1.2. Heuristics
For simplicity, in this subsection we focus on the case of hard-spheres i.e. $\gamma=1$ and $b=1$, which is the prototypical hard potential model. We write the weighted (1.13) as
${\langle}v{\rangle}^{k}Q^{+}(f,g)(v)=2\int_{\mathbb{R}^{3}\times{\mathbb{S}}^{2}}\frac{|u|{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{k}}f_{k}(v^{*})g_{k}(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma dv_{1}.$
The central object in our analysis is the ratio
(1.15) $\mathcal{R}:=\frac{|u|{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{k}},\quad u=v-v_{1},\quad(v,v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}.$
We distinguish two types of collisions: energy-distributing and energy-absorbing. Each of these requires a different argument.
Energy-distributing collisions
In most cases, energy is distributed among the outgoing particles, i.e. $|v^{*}|^{2}$ is comparable to the energy as well. We call these collisions energy-distributing. In Figure 1, we show an example of such a collision.
Figure 1. Energy-distributing collisions i.e. $|v^{*}|\approx|v_{1}^{*}|\approx E^{1/2}$
Now, since $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$, we can easily estimate
$\mathcal{R}\lesssim(1+E)^{\frac{1-k}{2}}\lesssim 1,$
as long as $k\geq 1$. In other words, in the most likely scenario, the weights offset the growth of the potential and the operator behaves like a Maxwell molecule. For this reason, we provide an exhaustive treatment of Maxwell molecules, which is of independent interest. In fact, we improve and extend existing results for such operators. As mentioned, convolution estimates without additional angular cut-offs for Maxwell molecules have been proved in [1, 2]. However, in these works the implicit constants in the estimates for $b=1$ are finite only when $r<3$. We opt for a self-contained, purely kinetic approach that allows us to extend the existing estimates on Maxwell molecules for $b$ satisfying (H.1)–(H.3) to arbitrarily large $r$ (even $r=\infty$), as long as $p<3$. This is the content of Theorem 3.1 ($r=\infty$) and Theorem 3.2 ($r<\infty$). We stress that our technique is also simpler, and only relies on elementary kinetic tools. Moreover, most of the aforementioned papers that introduce angular cut-offs in the gain operator fall into this category, since the cut-offs allow only energy-distributing collisions to occur. Our method could be used to revisit and extend these works as well.
Energy-absorbing collisions
The second, more pathological scenario consists of one of the outgoing particles accumulating almost all the energy of the collision i.e. $|v_{1}^{*}|\approx E^{1/2}$, $v^{*}\approx 0$. We call those energy-absorbing collisions. They distinguish hard potentials from Maxwell molecules. Following the billiard analogy, they can be visualized as two balls colliding, one staying still after the collision, while the other absorbs all of the energy. In Figures 2, 3, we provide examples of the two prototype energy-absorbing collisions, depending on whether $|v_{1}|\gtrsim|v|$ or $|v_{1}|\ll|v|$.
Figure 2. Energy-absorbing collision with $|v_{1}|\gtrsim|v|$, $v^{*}_{1}\approx v+v_{1}$
Figure 3. Energy-absorbing collisions with $|v_{1}|\ll|v|$, $v_{1}^{*}\approx v$
It is easy to see that energy-absorbing collisions happen with non-trivial probability. For example when $\sigma\approx\frac{2V}{|u|}$, the collisional law (1.3) yields $v^{*}\approx 0,\quad v_{1}^{*}\approx V$. However, in this case $\mathcal{R}\approx|u|=|v^{*}-v_{1}^{*}|\approx|v_{1}^{*}|\approx E^{1/2}$, where for the second equality we used (1.6). This leads to a growth of moments which cannot be controlled point-wise. Hence a crude bound would result in a moment-increasing estimate, which is not sufficient to understand the time evolution of the system, even locally in time.
However, despite the point-wise growth, we show that these singular energy-absorbing collisions do not drastically affect the averaging behavior of the gain operator. In other words, moment-preserving estimates still hold for the gain operator. The key idea is that, although energy-absorbing collisions occur with non-trivial probability, they are statistically rare. This is quantified by a cancellation and singular angular integration, see Lemma 2.5 (Cancellation Lemma) below. This estimate plays a crucial role in our analysis.
1.3. Main results
We now present the main results of this paper more formally. Given $k\geq 0$ and $1\leq p,q\leq\infty$, we define the space
(1.16) $\bm{\mathcal{X}_{p,q}^{k}}:=\left\{f\in{\langle}v{\rangle}^{-k}(L_{v}^{p}\cap L_{v}^{q})\,:\,\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}<\infty\right\},$
with norm
(1.17) $\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}:=\max\{\|{\langle}v{\rangle}^{k}f\|_{L^{p}_{v}},\|{\langle}v{\rangle}^{k}f\|_{L^{q}_{v}}\}.$
Note that the convergence of a sequence $\{f_{n}\}_{n=1}^{\infty}\subset\bm{\mathcal{X}_{p,q}^{k}}$ to some $f\in\bm{\mathcal{X}_{p,q}^{k}}$ can be equivalently characterized as follows:
[…]
It is straightforward to check that $(\bm{\mathcal{X}_{p,q}^{k}},\|\cdot\|_{\bm{\mathcal{X}_{p,q}^{k}}})$ is a Banach space.
We now state the main result of this paper:
Theorem 1.1.
Let $0<\gamma\leq 1$ and $b$ satisfying (H.1)–(H.3). Consider $1<p<2<q<\infty$, $2<r\leq\infty$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$, and $k\geq l>1$ with $k\geq 2+\gamma-\frac{2}{q}$, $l>2-\frac{2}{q}$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k-1+\gamma}L^{r}_{v}$ which satisfies the estimate
(1.19) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in\bm{\mathcal{X}_{p,q}^{k}}.$
Remark 1.2.
The first term on the right-hand side of (1.19) roughly corresponds to the piece of the operator where $|v_{1}|\ll|v|$ (see Figure 3), while the second term corresponds to $|v_{1}|\gtrsim|v|$, see Figure 2. The relative sizes of these velocities will be carefully quantified in terms of $|\widehat{\nu}\cdot\sigma|$, where $\nu=v-v_{1}^{*}$. As we will see, the second term controls the operator as long as $|\widehat{\nu}\cdot\sigma|>{\langle}v{\rangle}^{-1/2}$, while the first term arises in the singular region where $|\widehat{\nu}\cdot\sigma|<{\langle}v{\rangle}^{-1/2}$.
Remark 1.3.
In the case of a hard potential less singular than hard-spheres i.e. $\gamma<1$, our estimates are in fact moment-decreasing.
Remark 1.4.
Unlike for polynomial weights, we do not expect moment-preserving bounds on $Q^{+}$ to hold. Indeed, consider Gaussian weights of the form $e^{|v|^{2}}$. To prove the corresponding weighted estimates, one would need to control the ratio
$\mathcal{R}_{exp}:=\frac{|u|^{\gamma}e^{|v|^{2}}}{e^{|v^{*}|^{2}}e^{|v_{1}^{*}|^{2}}}=|u|^{\gamma}e^{|v|^{2}-|v^{*}|^{2}-|v_{1}^{*}|^{2}}=|u|^{\gamma}e^{-|v_{1}|^{2}},$
by conservation of energy. Now for $|v|\gg 1$ and $|v_{1}|\ll 1$, we have $\mathcal{R}_{exp}\approx|v|^{\gamma}$. However, unlike for polynomial weights, this quantity does not depend on the scattering angle $\sigma.$ Therefore collisional averaging cannot offset the growth induced by the potential.
Organization of the paper
The paper is organized as follows:
  • In Section 2, we provide the necessary kinetic tools that will be used in the rest of the paper. Namely, we introduce the Bobylev variables, prove Lemma 2.2 and Lemma 2.5, which are instrumental in our analysis.
  • In Section 3, we prove the convolution estimates for Maxwell molecules. Theorem 3.1 provides the convolution estimate for $r=\infty$, while Theorem 3.2 provides the convolution estimate for $r<\infty$.
  • In Section 4, we prove Theorem 1.1 which provides moment-preserving convolution estimates for hard-spheres and is the heart of our contribution.
Notation
Throughout this paper, given $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$, we will write
(1.20) $v^{*}=V-\frac{|u|}{2}\sigma,\ \ \ \ v_{1}^{*}=V+\frac{|u|}{2}\sigma,\ \ \ \ V=\frac{v+v_{1}}{2},\ \ \ u=v-v_{1},\ \ \ E=|v|^{2}+|v_{1}|^{2}.$
We will also use the following notation:
  • We write $C_{c}({\mathbb{R}}^{3})$ for the space of continuous, compactly supported functions in ${\mathbb{R}}^{3}$, $C_{c}^{\infty}({\mathbb{R}}^{3})$ for the space of smooth, compactly supported functions and $C_{b}({\mathbb{R}}^{3})$ for the space of continuous and bounded functions.
  • We write $A\lesssim B,$ to mean that there exists a numerical constant $C>0$, independent of $A,B$, such that $A\leq CB.$
  • We write $A\approx B$ to signify that both $A\lesssim B$ and $B\lesssim A$ hold.
  • We use the standard Japanese bracket notation: $\langle v\rangle:=\sqrt{1+|v|^{2}},$ where $|\cdot|$ denotes the $\ell_{v}^{2}$–norm of the vector $v\in\mathbb{R}^{3}.$
  • For $v\in\mathbb{R}^{3}$ with $v\neq 0$, we denote $\widehat{v}:=v/|v|.$
Acknowledgments
I.A. was supported by NSF grant No. DMS-2418020 and the PSC-CUNY Research Award 68653-00 56. The authors are also grateful to the Institute for Advanced Study for providing a peaceful environment to conduct this research during July 2025.
2. Technical toolbox
In this section, we present some technical results fundamental to our analysis. First, we introduce the so-called Bobylev variables [9, 10], and then employ them to prove key angular averaging estimates. Namely, Lemma 2.2 will be used to treat Maxwell molecules, while Lemma 2.5 captures the cancellation property of the gain collisional operator, which allows us to control energy-absorbing collisions and treat hard potentials. Finally, we record the well-known pre-post collisional change of variables in Lemma 2.6.
2.1. Bobylev variables
Given $\sigma\in{\mathbb{S}}^{2}$, we define the maps $R_{\sigma}^{+},R_{\sigma}^{-}:{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ by
(2.1) $R_{\sigma}^{+}(y)$ $=\frac{y}{2}+\frac{|y|}{2}\sigma,$
(2.2) $R_{\sigma}^{-}(y)$ $=\frac{y}{2}-\frac{|y|}{2}\sigma.$
One can readily verify that the following relations hold for any $y\in{\mathbb{R}}^{3}$:
(2.3) $R_{\sigma}^{+}(y)+R_{\sigma}^{-}(y)$ $=y,$
(2.4) $R_{\sigma}^{+}(y)\cdot R_{\sigma}^{-}(y)$ $=0,$
(2.5) $|R_{\sigma}^{+}(y)|^{2}+|R_{\sigma}^{-}(y)|^{2}$ $=|y|^{2}.$
Note that given $v,v_{1}\in{\mathbb{R}}^{3}$, $\sigma\in{\mathbb{S}}^{2}$, we can connect $R_{\sigma}^{+},R_{\sigma}^{-}$ with $v^{*},v_{1}^{*}$ as follows:
(2.6) $v^{*}$ $=v-R_{\sigma}^{+}(u),$
(2.7) $v_{1}^{*}$ $=v-R_{\sigma}^{-}(u).$
We now state some finer substitution properties of $R_{\sigma}^{+},R_{\sigma}^{-}$ that will be important in quantifying the collisional averaging behavior of the gain operator. The proof of the following result can be found in e.g. the work of Arsenio [7], as well as our previous work [3]. For convenience of the reader we present a detailed proof.
Proposition 2.1.
Let $\sigma\in{\mathbb{S}}^{2}$ and $\epsilon\in\{+,-\}$. Then the map
$R_{\sigma}^{\epsilon}:\{y\in{\mathbb{R}}^{3}:y\cdot\sigma\neq-\epsilon|y|\}\to\{\nu\in{\mathbb{R}}^{3}:\epsilon\,(\nu\cdot\sigma)>0\},$
is a diffeomorphism with inverse
(2.8) $\big(R_{\sigma}^{\epsilon}\big)^{-1}(\nu)=2\nu-\frac{|\nu|}{(\widehat{\nu}\cdot\sigma)}\sigma,$
and Jacobian
(2.9) $\text{Jac}\,(R_{\sigma}^{\epsilon})^{-1}(\nu)=\frac{4}{(\widehat{\nu}\cdot\sigma)^{2}}.$
Moreover, for any $y\in{\mathbb{R}}^{3}$ with $y\cdot\sigma\neq-\epsilon|y|$, we have
(2.10) $|R_{\sigma}^{\epsilon}(y)\cdot\sigma|$ $=\epsilon\,(R_{\sigma}^{\epsilon}(y)\cdot\sigma),$
(2.11) $|y|$ $=\frac{|R_{\sigma}^{\epsilon}(y)|}{|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|},$
(2.12) $\widehat{y}\cdot\sigma$ $=\epsilon\left(2|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|^{2}-1\right).$
Finally for $y\in{\mathbb{R}}^{3}$ with $y\cdot\sigma\neq\pm|y|$, we have
(2.13) $|\widehat{R}_{\sigma}^{+}(y)|^{2}+|\widehat{R}_{\sigma}^{-}(y)|^{2}=1.$
Proof.
Let $\sigma\in{\mathbb{S}}^{2}$ and $\epsilon\in\{+,-\}$. Let us also denote
$A_{\sigma}^{\epsilon}$ $=\{y\in{\mathbb{R}}^{3},\>\,y\cdot\sigma\neq-\epsilon|y|\},$
$B_{\sigma}^{\epsilon}$ $=\{\nu\in{\mathbb{R}}^{3}\,:\,\epsilon\,(\nu\cdot\sigma)>0\}.$
First, for $y\in A_{\sigma}^{\epsilon}$ we clearly have $R_{\sigma}^{\epsilon}(y)\cdot\sigma\neq 0$. Moreover, since $\epsilon\in\{+,-\}$, we obtain
(2.14) $\begin{split}\epsilon(R_{\sigma}^{\epsilon}(y)\cdot\sigma)&=\epsilon\left(\frac{y\cdot\sigma}{2}+\epsilon\frac{|y|}{2}\right)=\frac{|y|}{2}(1+\epsilon\,\widehat{y}\cdot\sigma)\\ &=\frac{|y|}{2}|1+\epsilon\,\widehat{y}\cdot\sigma|=\frac{|y|}{2}|\epsilon+\widehat{y}\cdot\sigma|\\ &=\left|\frac{y\cdot\sigma}{2}+\epsilon\frac{|y|}{2}\right|=|R_{\sigma}^{\epsilon}(y)\cdot\sigma|,\end{split}$
thus $R_{\sigma}^{\epsilon}:A_{\sigma}^{\epsilon}\to B_{\sigma}^{\epsilon}$, and moreover (2.10) holds.
In order to show that $R_{\sigma}^{\epsilon}$ is a diffeomorphism with inverse given by (2.8), consider $y\in A_{\sigma}^{\epsilon}$, $\nu\in B_{\sigma}^{\epsilon}$ with $\nu=R_{\sigma}^{\epsilon}(y)$. Then $y=2\nu-\epsilon|y|\sigma$, thus
$|y|^{2}=4|\nu|^{2}-4\epsilon|y|(\nu\cdot\sigma)+|y|^{2}\Rightarrow|y|=\frac{|\nu|}{\epsilon(\widehat{\nu}\cdot\sigma)},$
and (2.8) follows. Moreover, it is clear that $R_{\sigma}^{\epsilon},(R_{\sigma}^{\epsilon})^{-1}$ are continuously differentiable in $A_{\sigma}^{\epsilon},B_{\sigma}^{\epsilon}$ respectively, thus (2.8) follows. Moreover, it is clear that $R_{\sigma}^{\epsilon},\,(R_{\sigma}^{\epsilon})^{-1}$ are continuously differentiable in $A_{\sigma}^{\epsilon},B_{\sigma}^{\epsilon}$, thus $R_{\sigma}^{\epsilon}:A_{\sigma}^{\epsilon}\to B_{\sigma}^{\epsilon}$ is a diffeomorphism.
To compute the Jacobian, we differentiate (2.8) in $\nu$, to obtain
$D(R_{\sigma})^{-1}(\nu)=2I_{3}-\sigma\nabla^{T}f(\nu),$
where $I_{3}$ is the identity matrix and $f(\nu)=\frac{|\nu|^{2}}{(\nu\cdot\sigma)}$. Thus, we have111we use the Linear Algebra identity $\det(\lambda I_{n}+vw^{T})=\lambda^{n}(1+\lambda^{-1}v\cdot w)$, where $v,w\in{\mathbb{R}}^{n}$, $n\in\mathbb{N}$ and $\lambda\neq 0$.
(2.15) $\text{Jac}(R_{\sigma})^{-1}(\nu)=\det\left(2I_{3}-\sigma\nabla^{T}f(\nu)\right)=8\left(1-\frac{1}{2}\nabla f(\nu)\cdot\sigma\right).$
We readily compute
$\nabla f(\nu)\cdot\sigma=\frac{2(\nu\cdot\sigma)^{2}-|\nu|^{2}}{(\nu\cdot\sigma)^{2}}=2-\frac{|\nu|^{2}}{(\nu\cdot\sigma)^{2}},$
which, combined with (2.15), gives (2.9).
It remains to prove (2.11)–(2.13). Indeed, since $y=2R_{\sigma}^{\epsilon}(y)-\epsilon|y|\sigma$, we have
$|y|^{2}=\Big|2R_{\sigma}^{\epsilon}(y)-\epsilon|y|\sigma\Big|^{2}=4|R_{\sigma}^{\epsilon}(y)|^{2}+|y|^{2}-4\epsilon|y|(R_{\sigma}^{\epsilon}(y)\cdot\sigma),$
thus using (2.10) we obtain
$|R_{\sigma}^{\epsilon}(y)|^{2}=\epsilon|y|(R_{\sigma}^{\epsilon}(y)\cdot\sigma)=|y|\,|R_{\sigma}^{\epsilon}(y)\cdot\sigma|=|y|\,|R_{\sigma}^{\epsilon}(y)|\,|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|,$
and (2.11) follows.
Now using (2.11), (2.10), we obtain
$y\cdot\sigma$ $=2(R_{\sigma}^{\epsilon}(y)\cdot\sigma)-\epsilon|y|=\epsilon|y|\left(2\frac{\epsilon(R_{\sigma}^{\epsilon}(y)\cdot\sigma)}{|y|}-1\right)$
$=\epsilon|u|\left(2\frac{|R_{\sigma}^{\epsilon}(y)\cdot\sigma|}{|y|}-1\right)=\epsilon|y|\left(2|\widehat{R}_{\sigma}^{\epsilon}(y)\cdot\sigma|^{2}-1\right),$
and (2.12) follows.
Finally, for $y\in{\mathbb{R}}^{3}$ with $y\cdot\sigma\neq\pm|y|$ we obtain (2.13) by combining (2.5) with (2.11). ∎
2.2. Basic collisional averaging
Here, we use Lemma 2.1 to obtain the main averaging estimates when moments do not increase. Those estimates will be particularly useful for understanding Maxwell molecules, see Section 3.
Lemma 2.2.
Let $\beta_{1}<1$, $\beta_{2}>1/2$ and $1\leq p<\infty$. For $f\in C_{c}({\mathbb{R}}^{3})$, we denote
(2.16) $\mathcal{I}^{1}_{\beta_{1},p}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\mathds{1}_{\widehat{u}\cdot\sigma>0}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{1}}}\,d\sigma\,dv_{1},$
(2.17) $\mathcal{I}^{2}_{\beta_{2},p}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v_{1}^{*})|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{2}}\,d\sigma\,dv_{1}.$
Then, there hold the estimates
(2.18) $\|\mathcal{I}_{\beta_{1},p}^{1}[f]\|_{L^{\infty}_{v}}$ $\lesssim\|f\|_{L^{p}_{v}}^{p},$
(2.19) $\|\mathcal{I}_{\beta_{2},p}^{2}[f]\|_{L^{\infty}_{v}}$ $\lesssim\|f\|_{L^{p}_{v}}^{p}.$
Proof.
Consider $f\in C_{c}({\mathbb{R}}^{3})$. We first prove (2.18). Let us write $\chi(x):=\mathds{1}_{(0,\infty)}(x)$ and fix $v\in{\mathbb{R}}^{3}$. Using (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ we write
$\mathcal{I}_{\beta_{1},p}^{1}[f](v)$ $=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(u))|^{p}\chi(2|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{\beta_{1}}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(y))|^{p}\chi(2|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}-1)}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{\beta_{1}}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ and obtain
$\mathcal{I}_{\beta_{1},p}^{1}[f](v)$ $\approx\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-\nu)|^{p}\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{\beta_{1}}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\int_{{\mathbb{S}}^{2}}\frac{\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{\beta_{1}}}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{0}^{1}\frac{\chi(2x^{2}-1)}{x^{2}(1-x^{2})^{\beta_{1}}}\,dx\right)\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{1/\sqrt{2}}^{1}\frac{1}{(1-x^{2})^{\beta_{1}}}\,dx\right)\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{1/\sqrt{2}}^{1}\frac{1}{(1-x)^{\beta_{1}}}\,dx\right)\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\,d\nu=\|f\|_{L^{p}_{v}}^{p},$
where we used the fact that $\beta_{1}<1$ for the convergence of the integral in $x$ and translation invariance for the last equality. Estimate (2.18) is proved.
We now prove (2.19). Fix again $v\in{\mathbb{R}}^{3}$. Using (2.7) and (2.12), we write
$\mathcal{I}_{\beta_{2},p}^{2}[f](v)$ $=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f(v-R_{\sigma}^{-}(u))|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{2}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f(v-R_{\sigma}^{-}(y))|^{p}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|^{2\beta_{2}}\,dy\,d\sigma.$
Using again Proposition 2.1 to substitute $\nu:=R_{\sigma}^{-}(y)$, we obtain
$\mathcal{I}_{\beta_{2},p}^{2}[f](v)$ $\approx\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}|\widehat{\nu}\cdot\sigma|^{2(\beta_{2}-1)}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\int_{{\mathbb{S}}^{2}}|\widehat{\nu}\cdot\sigma|^{2(\beta_{2}-1)}\,d\sigma\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\left(\int_{0}^{1}x^{2(\beta_{2}-1)}\,dx\right)\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|f(v-\nu)|^{p}\,d\nu=\|f\|_{L^{p}_{v}}^{p},$
where we used the fact that $\beta_{2}>1/2$ for the convergence of the integral in $x$. Estimate (2.19) is proved. ∎
2.3. Cancellation Lemma
Now we use Lemma 2.1 to prove Lemma 2.5 which encodes the cancellation mechanism of the gain collisional operator for hard potentials. First, we present an important point-wise cancellation estimate in Lemma 2.3 and an elementary singular integral estimate in Lemma 2.4:
Lemma 2.3.
For any $v,v_{1}\in{\mathbb{R}}^{3}$, $\sigma\in{\mathbb{S}}^{2}$, the following bounds hold
(2.20) ${\langle}v^{*}{\rangle}$ $\geq{\langle}v_{1}^{*}{\rangle}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)^{1/2},\quad\lambda_{v_{1}^{*}}=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}},$
(2.21) ${\langle}v_{1}^{*}{\rangle}$ $\geq{\langle}v^{*}{\rangle}\left(1-\lambda_{v^{*}}|\widehat{v^{*}}\cdot\sigma|\right)^{1/2},\quad\lambda_{v^{*}}=\frac{|v^{*}|^{2}}{{\langle}v^{*}{\rangle}^{2}}.$
Proof.
Fix $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$. By (1.3) we have $v^{*}=v_{1}^{*}-|u|\sigma$, so we compute
$|v^{*}|^{2}=|v_{1}^{*}|^{2}+|u|^{2}-2|u||v_{1}^{*}|(\widehat{v}_{1}^{*}\cdot\sigma)\geq(|v_{1}^{*}|^{2}+|u|^{2})\left(1-\frac{2|u||v_{1}^{*}|}{|v_{1}^{*}|^{2}+|u|^{2}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)\geq|v_{1}^{*}|^{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|),$
where we used the basic inequality $2|u||v_{1}^{*}|\leq|u|^{2}+|v_{1}^{*}|^{2}$. Therefore
${\langle}v_{1}^{*}{\rangle}^{2}=1+|v_{1}^{*}|^{2}\geq(1+|v_{1}^{*}|^{2})\left(1-\frac{|v_{1}^{*}|^{2}}{1+|v_{1}^{*}|^{2}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)={\langle}v_{1}^{*}{\rangle}^{2}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|\right),$
where $\lambda_{v_{1}^{*}}=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}}$. Estimate (2.21) is proved identically, using the expression $v_{1}^{*}=v^{*}+|u|\sigma$ instead. ∎
Lemma 2.4.
Let $l>2$. Then for any $\widehat{\eta}\in{\mathbb{S}}^{2}$ and any $0<\lambda<1$, the following bound holds
(2.22) $\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|)^{l/2}}\,d\sigma\lesssim\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2}.$
Proof.
Let $l>2$ and fix $\eta\in{\mathbb{S}}^{2}$ as well as $0<\lambda<1$. Integrating in spherical coordinates, we obtain
$\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|)^{l/2}}\,d\sigma$ $\approx\int_{0}^{1}\frac{1}{(1-\lambda x)^{l/2}}\,dx=\frac{1}{\lambda}\int_{0}^{\lambda}\frac{1}{(1-z)^{l/2}}\,dz\leq\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2},$
since $l>2$. ∎
We are now in the position to prove the main cancellation estimate:
Lemma 2.5 (Cancellation lemma).
Let $l>2$ and $1\leq p<\infty$. For $g\in C_{c}({\mathbb{R}}^{3})$, we denote
(2.23) $\mathcal{J}_{l,p}^{0}[g](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{+}(u)|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|g(v^{*})|^{p}\mathds{1}_{|v^{*}|>2}\,d\sigma\,dv_{1}$
(2.24) $\mathcal{J}_{l,p}^{1}[g](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)|^{2}}{{\langle}v^{*}{\rangle}^{l}}|g(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}.$
Then there hold the estimates
(2.25) $\|\mathcal{J}_{l,p}^{0}[g]\|_{L_{v}^{\infty}}$ $\lesssim\|g\|_{L_{v}^{p}}^{p}.$
(2.26) $\|\mathcal{J}_{l,p}^{1}[g]\|_{L_{v}^{\infty}}$ $\lesssim\|g\|_{L_{v}^{p}}^{p}.$
Proof.
We prove estimate (2.26) as estimate (2.25) follows in an entirely analogous way. Throughout the proof, given $a\in{\mathbb{R}}^{3}$ with $a\neq 0$, we will denote $\lambda_{a}:=\frac{|a|^{2}}{{\langle}a{\rangle}^{2}}\in(0,1)$. Consider $g\in C_{c}({\mathbb{R}}^{3})$ and fix $v\in{\mathbb{R}}^{3}$. We use (2.20), (2.7) and the substitution $y:=u=v-v_{1}$ to bound
$\mathcal{J}_{l,p}^{1}[g](v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|}{{\langle}v_{1}^{*}{\rangle}^{l}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)^{l/2}}|g(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|g(v-R_{\sigma}^{-}(u))|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{\left\langle v-R_{\sigma}^{-}(u)\right\rangle^{l-2}\left(1-\lambda_{v-R_{\sigma}^{+}(u)}|\widehat{(v-R_{\sigma}^{-}(u))}\cdot\sigma|\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(u)|>2}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|g(v-R_{\sigma}^{-}(y))|^{p}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|}{{\langle}v-R_{\sigma}^{-}(y){\rangle}^{l-2}\left(1-\lambda_{v-R_{\sigma}^{-}(y)}|\widehat{(v-R_{\sigma}^{-}(y))}\cdot\sigma|\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(y)|^{2}>2}\,dy\,d\sigma.$
Using Proposition 2.1 to substitute $\nu:=R_{\sigma}^{-}(y)$, we obtain
$\mathcal{J}_{l,p}^{1}[g](v)\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{2-l}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|\right)^{l/2}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{2-l}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\int_{{\mathbb{S}}^{2}}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|\right)^{l/2}}\,d\sigma\,d\nu.$
Then, since $l>2$ and $\lambda_{v-\nu}=\frac{|v-\nu|^{2}}{{\langle}v-\nu{\rangle}^{2}}\in(0,1)$, Lemma 2.4 implies
$\mathcal{J}_{l,p}^{1}[g](v)$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{2-l}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\frac{(1-\lambda_{v-\nu})^{1-l/2}}{\lambda_{v-\nu}}\,d\nu$
$\lesssim\int_{{\mathbb{R}}^{3}}\frac{{\langle}v-\nu{\rangle}^{2}}{|v-\nu|^{2}}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\,d\nu$
$\approx\int_{{\mathbb{R}}^{3}}|g(v-\nu)|^{p}\,d\nu=\|g\|_{L_{v}^{p}}^{p}.$
Estimate (2.26) is proved. ∎
2.4. Pre-post collisional change of variables
Finally, we record the well-known pre-post collisional change of variables, see e.g. [37], which will be useful in the $L^{r}$– estimates. We provide the proof for convenience of the reader.
Lemma 2.6.
The map $T:(v,v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\to(v^{*},v_{1}^{*},\eta)\in{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}$ where
(2.27) $\begin{cases}v^{*}=\frac{v+v_{1}}{2}-\frac{|v-v_{1}|}{2}\sigma\\ v_{1}^{*}=\frac{v+v_{1}}{2}+\frac{|v-v_{1}|}{2}\sigma\\ \eta=-\widehat{u}=\frac{v_{1}-v}{|v-v_{1}|}\end{cases}$
is an involution of ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}$. Moreover, for any non-negative and continuously differentiable function $F:{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times[-1,1]\to{\mathbb{R}}_{+}$, there holds the change of variables formula
(2.28) $\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v^{*},v_{1}^{*},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v,v_{1},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv.$
Remark 2.7.
It is customary in literature to abbreviate notation and write $(v,v_{1})\to(v^{*},v_{1}^{*})$ to indicate the above change of variables, however we stress that it is not correct to perform this change of variables for a fixed $\sigma\in{\mathbb{S}}^{2}$.
Proof.
Writing $T^{2}:=T\circ T$, we compute
$T^{2}(v,v_{1},\sigma)$ $=T(v^{*},v_{1}^{*},\eta)=\begin{pmatrix}\frac{v^{*}+v_{1}^{*}}{2}-\frac{|v^{*}-v_{1}^{*}|}{2}\eta\\ \frac{v^{*}+v_{1}^{*}}{2}+\frac{|v^{*}-v_{1}^{*}|}{2}\eta\\ \frac{v_{1}^{*}-v^{*}}{|v^{*}-v_{1}^{*}|}\end{pmatrix}=\begin{pmatrix}\frac{v+v_{1}}{2}-\frac{|v-v_{1}|}{2}\eta\\ \frac{v+v_{1}}{2}+\frac{|v-v_{1}|}{2}\eta\\ \frac{v_{1}^{*}-v^{*}}{|v-v_{1}|}\end{pmatrix}=\begin{pmatrix}v\\ v_{1}\\ \sigma\end{pmatrix}$.
Therefore $T^{2}=Id$, so $T$ is an involution.
Second, writing $u^{*}:=v^{*}-v_{1}^{*}=-|u|\sigma$, we compute
$u^{*}\cdot\eta=-|u|\sigma\cdot(-\widehat{u})=u\cdot\sigma,$
thus by (1.6) we obtain $\widehat{u}\cdot\sigma=\widehat{u^{*}}\cdot\eta.$ Therefore
$\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v^{*},v_{1}^{*},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v^{*},v_{1}^{*},\widehat{u^{*}}\cdot\eta)\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}F(v,v_{1},\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv,$
since $T$ is an involution and thus has unitary Jacobian. ∎
3. Convolution estimates for Maxwell molecules
In the section, we prove the convolution estimates for Maxwell molecules. Given $b$ satisfying (H.1)–(H.3), we define the Maxwell bilinear collisional operator with angular cross-section $b$ as
(3.1) $Q_{M,b}^{+}(f,g)=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\mathds{1}_{\widehat{u}\cdot\sigma>0}\,d\sigma\,dv_{1}.$
We note that, when $f=g$, (H.2) implies
$Q_{M,b}^{+}(f,f)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\,d\sigma\,dv_{1},$
thus (3.1) extends (1.9) for $\gamma=0$. Slightly abbreviating notation we will be writing $Q^{+}_{M}:=Q^{+}_{M,1}$ for the Maxwell operator with constant cross-section equal to $1$.
As mentioned, when $b$ is constant, the previous works [1, 2] only provide an estimate when $r<3$. Here, we are able to extend the estimates to arbitrary $r$, as long as $p<3$.
3.1. The $L_{v}^{\infty}$– estimate
First, we prove the convolution estimates corresponding to $r=\infty$.
Theorem 3.1.
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then, $Q_{M,b}^{+}$ extends uniquely to a bounded, bilinear operator $Q_{M,b}^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{\infty}_{v}$ satisfying the estimate
(3.2) $\|Q^{+}_{M,b}\|_{L_{v}^{\infty}}\lesssim\|f\|_{L_{v}^{p}}\|g\|_{L_{v}^{q}}.$
For $p=1$, $q=\infty$, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:L^{1}_{v}\times C_{b}({\mathbb{R}}^{3})\to L^{\infty}_{v}$ satisfying the estimate
(3.3) $\|Q^{+}_{M,b}\|_{L^{\infty}_{v}}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}}.$
Proof.
We start by proving the first claim. By density, it suffices to show (3.2) for $f,g\in C_{c}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}({\mathbb{R}}^{3})$. Since $p<3$, we have $\frac{1}{2q}<\frac{1}{p}$. Hence we can choose $\alpha>1/2$ such that $\alpha p<q$. We use (H.3) and Hölder’s inequality to bound
$|Q_{M,b}^{+}(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f(v^{*})|\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/q}}\right)\left(|g(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/q}\right)\,d\sigma\,dv_{1}$
(3.4) $\leq I_{1}^{1/p}(v)I_{2}^{1/q}(v),$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha p/q}}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g(v_{1}^{*})|^{q}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}\,d\sigma\,dv_{1}.$
Since $\alpha p<q$, $\alpha>1/2$, Lemma 2.2 implies $I_{1}(v)\lesssim\|f\|_{L^{p}_{v}}^{p}$ and $I_{2}(v)\lesssim\|g\|_{L_{v}^{q}}^{q}$. Combining (3.4) with the bounds for $I_{1},I_{2}$, (3.2) follows. The claim is proved.
Now for the case $p=1$, $q=\infty$, again by density, it suffices to show (3.3) for $f\in C_{c}({\mathbb{R}}^{3})$. Assume $f\in C_{c}({\mathbb{R}}^{3})$ and $g\in C_{b}({\mathbb{R}}^{3})$. By (H.3) and (2.18), we bound
(3.5) $|Q^{+}_{M,b}(f,g)|$ $\lesssim\|g\|_{L^{\infty}_{v}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\lesssim\|f\|_{L^{1}_{v}}\|g\|_{L^{\infty}_{v}},$
and (3.3) follows. The claim is proved.
3.2. The $L_{v}^{r}$– estimate
Now, we prove the convolution estimates for $1\leq r<\infty$.
Theorem 3.2.
Let $b$ satisfying (H.1)–(H.3) and consider $1\leq p,q,r<\infty$ with $p<3$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$. Then, $Q_{M,b}^{+}$ extends uniquely to a bounded, bilinear operator $Q_{M,b}^{+}:L^{p}_{v}\times L^{q}_{v}\to L^{r}_{v}$ satisfying the estimate
(3.6) $\|Q_{M,b}^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f\|_{L_{v}^{p}}\|g\|_{L_{v}^{q}}.$
Proof.
By density, it suffices to prove (3.6) for $f,g\in C_{c}({\mathbb{R}}^{3})$. Fix $f,g\in C_{c}({\mathbb{R}}^{3})$. We treat the cases $1<p<3$ and $p=1$ separately.
$\bullet$ Case $1<p<3$ : Since $p<3$, we have $p^{\prime}>3/2$, thus
(3.7) $\frac{1}{2p^{\prime}}=\frac{1}{2}-\frac{1}{2q^{\prime}}-\frac{1}{2r}=\frac{1}{q^{\prime}}+\frac{1}{r}+\frac{1}{2}-\frac{3}{2q^{\prime}}-\frac{3}{2r}=\frac{1}{q^{\prime}}+\frac{1}{r}+\frac{3}{2p^{\prime}}-1<\frac{1}{q^{\prime}}+\frac{1}{r}.$
Hence we can choose $\alpha>1/2$ such that $\frac{\alpha}{p^{\prime}}<\frac{1}{q^{\prime}}+\frac{1}{r}$. Equivalently, we can write this condition as $1-\frac{p^{\prime}}{\alpha q^{\prime}}<\frac{p^{\prime}}{\alpha r}$. Pick $s\in{\mathbb{R}}$ with $1-\frac{p^{\prime}}{\alpha q^{\prime}}<s<\frac{p^{\prime}}{\alpha r}$. Then $\frac{(1-s)\alpha q^{\prime}}{p^{\prime}}<1$ and $\frac{s\alpha r}{p^{\prime}}<1$.
Fix $v\in{\mathbb{R}}^{3}$. Using (H.3) and Hölder’s inequality, we bound
$|Q_{M,b}^{+}(f,g)(v)|$ $\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f(v^{*})|^{p/q^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha(1-s)/p^{\prime}}}\right)\left(|g(v_{1}^{*})|^{q/p^{\prime}}|\widehat{R}^{-}_{\sigma}(u)\cdot\sigma|^{2\alpha/p^{\prime}}\right)$
$\hskip 113.81102pt\times\left(\frac{|f(v^{*})|^{p/r}|g(v_{1}^{*})|^{q/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha s/p^{\prime}}}\right)\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(3.8) $\lesssim I_{1}^{1/q^{\prime}}(v)I_{2}^{1/p^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|g(v_{1}^{*})|^{q}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha sr/p^{\prime}}}\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha(1-s)q^{\prime}/p^{\prime}}}\,d\sigma\,dv_{1}\,dv,$
$I_{2}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g(v_{1}^{*})|^{q}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}\,d\sigma\,dv_{1}\,dv.$
Since $\frac{(1-s)\alpha q^{\prime}}{p^{\prime}}<1$ and $\alpha>1/2$, Lemma 2.2 implies $I_{1}(v)\lesssim\|f\|_{L^{p}_{v}}^{p}$ and $I_{2}(v)\lesssim\|g\|_{L^{q}}^{q}$.
Combining (3.8) with the bounds for $I_{1},I_{2}$, we obtain
$|Q_{M,b}^{+}(f,g)(v)|\lesssim\|f\|_{L_{v}^{p}}^{p/q^{\prime}}\|g\|_{L_{v}^{q}}^{q/p^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|g(v_{1}^{*})|^{q}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha sr/p^{\prime}}}\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above bound to the $r$-th power, integrating in $v$, and using (2.28), we obtain
$\|Q_{M,b}^{+}(f,g)\|_{L_{v}^{r}}^{r}$ $\lesssim\|f\|_{L_{v}^{p}}^{r-p}\|g\|_{L_{v}^{q}}^{r-q}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}|g(v_{1}^{*})|^{q}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha sr/p^{\prime}}}\,d\sigma\,dv_{1}\,dv$
$=\|f\|_{L_{v}^{p}}^{r-p}\|g\|_{L_{v}^{q}}^{r-q}\int_{{\mathbb{R}}^{6}}|f(v)|^{p}|g(v_{1})|^{q}\left(\int_{{\mathbb{S}}^{2}}\frac{\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)|^{2\alpha sr/p^{\prime}}}\,d\sigma\right)\,dv_{1}\,dv.$
To estimate the inner sphere integral we use (2.12) to write
$\int_{{\mathbb{S}}^{2}}\frac{\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)|^{2\alpha sr/p^{\prime}}}\,d\sigma\approx\int_{{\mathbb{S}}^{2}}\frac{\chi(\widehat{u}\cdot\sigma)}{(1-\widehat{u}\cdot\sigma)^{s\alpha r/p^{\prime}}}\,d\sigma\approx\int_{0}^{1}\frac{1}{(1-x)^{s\alpha r/p^{\prime}}}\,dx\approx 1,$
since $\frac{s\alpha r}{p^{\prime}}<1$.
We conclude
(3.9) $\|Q_{M,b}^{+}(f,g)\|_{L^{r}_{v}}^{r}\lesssim\|f\|_{L_{v}^{p}}^{r-p}\|g\|_{L_{v}^{q}}^{r-q}\int_{{\mathbb{R}}^{6}}|f(v)|^{p}|g(v_{1})|^{q}\,dv_{1}\,dv=\|f\|_{L_{v}^{p}}^{r}\|g\|_{L_{v}^{q}}^{r}.$
$\bullet$ Case $p=1$ : We use (H.3) and Hölder’s inequality to bound
$|Q_{M,b}^{+}(f,g)(v)|$ $\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|^{1/r^{\prime}}\left(|f(v^{*})|^{1/r}|g(v_{1}^{*})|\right)\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(3.10) $\leq I_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
By (2.18), we have $I_{1}\lesssim\|f\|_{L^{1}_{v}}$. Combining (3.10) with the estimate on $I_{1}$, we obtain
$|Q_{M,b}^{+}(f,g)(v)|\lesssim\|f\|_{L^{1}_{v}}^{1/r^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the expression above to the $r$-th power and integrating in $v$, we obtain
(3.11) $\|Q_{M,b}^{+}(f,g)(v)\|_{L^{r}_{v}}^{r}$ $\lesssim\|f\|_{L^{1}_{v}}^{r-1}\left(\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f(v^{*})||g(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\,dv\right)\lesssim\|f\|_{L^{1}_{v}}^{r}\|g\|_{L^{r}_{v}}^{r}.$
Combining (3.9), (3.11), estimate (3.6) follows. The claim is proved. ∎
4. Convolution estimates for hard-spheres
In this final section, we prove our main result, Theorem 1.1, for the bilinear operator $Q^{+}(f,g)$ given in (1.13). We first establish the necessary estimates for smooth, compactly supported functions for $r=\infty$ and $2<r<\infty$ respectively.
4.1. The $L_{v}^{\infty}$– estimate
Proposition 4.1.
Let $0<\gamma\leq 1$ and $b$ satisfying (H.1)–(H.3). Consider $1<p<q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$ and $k\geq l>1$ with $k\geq\gamma+\frac{2}{p}$, $l>\frac{2}{p}$. Then, there holds the estimate
(4.1) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in C_{c}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>1$ with $k\geq\gamma+\frac{2}{p}$ and $l>2/p$. Fix $f,g\in C_{c}({\mathbb{R}}^{3})$ and denote $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$.
We use the triangle inequality to bound $Q^{+}$:
(4.2) $|Q^{+}(f,g)|\lesssim|Q^{sm}(f,g)|+|Q^{r}(f,g)|+|Q^{s_{1}}(f,g)|+|Q_{1}^{s_{2}}(f,g)|,$
where
$Q_{sm}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|\leq 4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1},$
$Q_{r}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{|v^{*}|\geq|v|/2}\,d\sigma\,dv_{1},$
$Q_{s,1}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
$Q_{s,2}^{+}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1},$
and given $v\in{\mathbb{R}}^{3}$, we denoted
(4.3) $\mathcal{B}_{1}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v|}{2},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|\geq{\langle}v{\rangle}^{-1/2}\right\},$
(4.4) $\mathcal{B}_{2}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v|}{2},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<{\langle}v{\rangle}^{-1/2}\right\}.$
Estimate for $Q_{sm}^{+}$
Using the inequality $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we bound
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k+1-\gamma}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v|\leq 4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim E^{\gamma/2}(1+E)^{-k/2}\lesssim 1,$
since $k>1$. Using the above estimate, the substitution $\sigma\to-\sigma$ for the second summand and recalling the Maxwell operator notation (3.1), we obtain
${\langle}v{\rangle}^{k+1-\gamma}|Q_{sm}^{+}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(-\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v_{1}^{*})g_{k}(v^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\approx Q_{M,1}^{+}(|f_{l}|,|g_{k}|)+Q_{M,1}^{+}(|g_{k}|,|f_{l}|).$
Then, since $p<2$, Theorem 3.1 implies
(4.5) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{sm}^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L^{q}_{v}}\|g_{k}\|_{L^{p}_{v}}.$
Estimate for $Q_{r}^{+}$
Using the inequality $|u|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we bound
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k+1-\gamma}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v|>4}\mathds{1}_{|v^{*}|\geq|v|/2}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim\frac{E^{\gamma/2}{\langle}v{\rangle}^{k+1-\gamma-l}}{(1+E)^{k/2}}\leq(1+E)^{\frac{1-l}{2}}\lesssim 1$
since $l>1$. Arguing as in the previous case, we obtain
(4.6) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{r}^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
It remains to estimate the singular operators $Q^{+}_{s,1}$ and $Q^{+}_{s,2}$.
We start with a preliminary observation: for $|v^{*}|<|v|/2$, we use (2.11), to write
$|v|\leq|v-v^{*}|+|v^{*}|\leq|R_{\sigma}^{+}(u)|+\frac{|v|}{2}\Rightarrow|v|\leq 2|R_{\sigma}^{+}(u)|=2|u||\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|.$
Now, since $0<\gamma\leq 1$ and $|u|\lesssim E^{1/2}$, we obtain
(4.7) $\frac{|u|^{\gamma}{\langle}v{\rangle}^{k+1-\gamma}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{|v^{*}|<|v|/2}\lesssim\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{|v^{*}|<|v|/2}.$
Estimate for $Q_{s,1}^{+}$
Using (4.7), the exponents’ compatibility condition, and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k+1-\gamma}|Q_{s,1}^{+}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}^{\frac{1}{q}-\frac{1}{p}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p}}\right)\left(\frac{{\langle}v{\rangle}^{2/p}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}$
(4.8) $\leq I_{1}^{1/q}(v)I_{2}^{1/p}(v),$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-\frac{q}{p}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{q(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q/p}}|f_{l}(v^{*})|^{q}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp}}|g_{k}(v_{1}^{*})|^{p}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $l>2/p$, we use (2.26) to bound
$I_{2}(v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp}}|g_{k}(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}\lesssim\|g_{k}\|_{L^{p}_{v}}^{p}.$
In order to estimate $I_{1}$, we use the definition (4.3) of $\mathcal{B}_{1}(v)$, followed by (2.6), (2.13) and the substitution $y:=u=v-v_{1}$ to write
$I_{1}(v)$ $\leq{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(u))|^{q}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{q(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q/p}}\,dv_{1}\,d\sigma$
$={\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(y))|^{q}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{q(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q/p}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound, as well as the fact that $q(k-\gamma)>p(k-\gamma)\geq 2$, to obtain
$I_{1}(v)$ $\lesssim{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\frac{|\widehat{\nu}\cdot\sigma|^{q(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q/p}}\,d\nu\,d\sigma$
$\leq{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q/p}}\,d\sigma\,d\nu$
$\approx{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q/p}}\,dx\,d\nu.$
Now, the inner integral can be estimated as follows
$\int_{0}^{1}$ $\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q/p}}\,dx\lesssim 1+\int_{1/\sqrt{2}}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q/p}}x\,dx\approx 1+\int_{{\langle}v{\rangle}^{-1}}^{1/2}z^{-q/p}\,dz\lesssim{\langle}v{\rangle}^{\frac{q}{p}-1},$
where we used the substitution $z:=1-x^{2}$, and the fact that $p<q$. We obtain
$I_{1}(v)\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\,d\nu=\|f_{l}\|_{L^{q}_{v}}^{q}.$
Combining (4.8) with the estimates for $I_{1},I_{2}$, we obtain
(4.9) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{s,1}^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
Estimate for $Q_{s,2}^{+}$
Using (4.7), the exponents’ compatibility condition, and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k+1-\gamma}|Q_{s,2}^{+}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}^{\frac{1}{p}-\frac{1}{q}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q}}\right)\left(\frac{{\langle}v{\rangle}^{2/q}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}$
(4.10) $\leq J_{1}^{1/p}(v)J_{2}^{1/q}(v),$
where
$J_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-\frac{p}{q}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p/q}}|f_{l}(v^{*})|^{p}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1},$
$J_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq}}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $ql>pl>2$, we use (2.26) to bound
$J_{2}(v)\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq}}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}\lesssim\|g_{k}\|_{L^{q}_{v}}^{q}.$
In order to estimate $J_{1}$, we use a symmetric argument as for $I_{1}$. Namely, we use the definition (4.4) of $\mathcal{B}_{2}(v)$, followed by (2.6), (2.13) and the substitution $y:=u=v-v_{1}$ to write
$J_{1}(v)$ $\leq{\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(u))|^{p}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q/p}}\,dv_{1}\,d\sigma$
$={\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(y))|^{p}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{p(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q/p}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound, as well as the fact that $p(k-\gamma)\geq 2$, to obtain
$J_{1}(v)$ $\lesssim{\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\mathds{1}_{|v|>4}\frac{|\widehat{\nu}\cdot\sigma|^{p(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{p/q}}\,d\nu\,d\sigma$
$\leq{\langle}v{\rangle}^{1-\frac{p}{q}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\mathds{1}_{|v|>4}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}<{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{p/q}}\,d\sigma\,d\nu$
$\approx{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p/q}}\,dx\,d\nu.$
The inner integral now can be estimated as follows
$\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p/q}}\,dx\lesssim\int_{0}^{1}\frac{\mathds{1}_{1-x<{\langle}v{\rangle}^{-1}}}{(1-x)^{p/q}}\approx\int_{0}^{{\langle}v{\rangle}^{-1}}z^{-p/q}\,dz\lesssim{\langle}v{\rangle}^{\frac{p}{q}-1},$
where we used the fact that $p<q$ for the convergence of the integral in $z$. We obtain
$J_{1}(v)\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\,d\nu=\|f_{l}\|_{L^{p}}^{p}.$
Combining (4.10) with the bounds for $J_{1},J_{2}$, we obtain
(4.11) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}_{s,2}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}.$
Conclusion of the proof
Combining (4.2) with (4.5), (4.6), (4.9), (4.11), estimate (4.1) follows.
4.2. The $L_{v}^{r}$– estimate
Proposition 4.2.
Let $0<\gamma\leq 1$ and $b$ satisfying (H.1)–(H.3). Consider $2<r<\infty$, $1<p<2<q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$ and $k\geq l>1$ with $k\geq 2+\gamma-\frac{2}{q}$, $l>2-\frac{2}{q}$. Then, there holds the estimate
(4.12) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in C_{c}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>1$ with $k\geq 2+\gamma-\frac{2}{q}$, $l>2-\frac{2}{q}$. Note that these assumptions imply $q^{\prime}(k-\gamma)\geq 2$ and $q^{\prime}l>2$. Consider $f,g\in C_{c}({\mathbb{R}}^{3})$, and denote again $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$.
We recall the decomposition (4.2) from Proposition 4.1.
Estimate for $Q_{sm}^{+}$, $Q_{r}^{+}$
The operators $Q_{sm}^{+},Q_{r}^{+}$ can be treated similarly to Proposition 4.1 using Theorem 3.2 instead of Theorem 3.1. Omitting the details, we obtain
(4.13) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{sm}^{+}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}},$
(4.14) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{r}^{+}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}+\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
We now estimate the singular operators $Q_{s,1}^{+}$ and $Q_{s,2}^{+}$.
Estimate for $Q_{s,1}^{+}$
Since $q>2$, we have $\frac{1}{r}<\frac{1}{p}-\frac{1}{q}=\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$. Hence we can choose $\alpha>1$ such that $\frac{\alpha}{r}<\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$. Relying on the exponents’ compatibility condition, we can write
(4.15) $1=\frac{1}{p^{\prime}}+\frac{1}{q^{\prime}}+\frac{1}{r}=\left(\frac{1}{p^{\prime}}-\frac{1}{q^{\prime}}+\frac{\alpha}{r}\right)+\frac{2}{q^{\prime}}+\frac{1-\alpha}{r}.$
We argue by duality; we fix $\phi\in L_{v}^{r^{\prime}}$, and estimate the quantity
$I(\phi):=\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{k+1-\gamma}Q_{s,1}^{+}(v)\phi(v)\,dv.$
Using (4.7), the fact that $|v|\lesssim|v_{1}^{*}|$, (4.15) and Hölder’s inequality, we bound
$|I(\phi)|$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}|\phi(v)|\,d\sigma\,dv_{1}\,dv$
$\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v_{1}^{*}{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}|\phi(v)|\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v_{1}^{*}{\rangle}^{\frac{1}{p^{\prime}}-\frac{1}{q^{\prime}}+\frac{\alpha}{r}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-1}|f_{l}(v^{*})|^{q/p^{\prime}}|\phi(v)|^{r^{\prime}/p^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}\right)$
$\hskip 85.35826pt\times\left(\frac{{\langle}v_{1}^{*}{\rangle}^{2/q^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q^{\prime}}|g_{k}(v_{1}^{*})|^{p/q^{\prime}}|\phi(v)|^{r^{\prime}/q^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left(\frac{{\langle}v_{1}^{*}{\rangle}^{\frac{1-\alpha}{r}}|f_{l}(v^{*})|^{q/r}|g_{k}(v_{1}^{*})|^{p/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/r}}\right)\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}}\,d\sigma\,dv_{1}\,dv$
(4.16) $\leq I_{1}^{1/p^{\prime}}I_{2}^{1/q^{\prime}}I_{3}^{1/r},$
where
$I_{1}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{1-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p^{\prime}(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}|f_{l}(v^{*})|^{q}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv,$
$I_{2}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq^{\prime}}}|g_{k}(v_{1}^{*})|^{p}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv,$
$I_{3}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v_{1}^{*}{\rangle}^{1-\alpha}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $q^{\prime}l>2$, we use (2.26) to bound
$I_{2}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq^{\prime}}}|g_{k}(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}\right)\,dv\lesssim\|g_{k}\|_{L^{p}_{v}}^{p}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
For $I_{1}$, since $|v|\lesssim|v_{1}^{*}|$ and $\frac{\alpha}{r}<\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$, we first bound
$I_{1}$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p^{\prime}(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}|f_{l}(v^{*})|^{q}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv.$
Then, since $p^{\prime}(k-\gamma)>q^{\prime}(k-\gamma)\geq 2$, we argue similarly as in the estimate of $I_{1}$ in the proof of Proposition 4.1, to obtain
$I_{1}$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{1-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}\,dx\,d\nu\,dv.$
The inner integral can be estimated as follows
$\int_{0}^{1}$ $\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}\,dx\lesssim 1+\int_{1/\sqrt{2}}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}x\,dx\approx 1+\int_{{\langle}v{\rangle}^{-1}}^{1/2}z^{-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}\,dz\lesssim{\langle}v{\rangle}^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})-1},$
where we used the substitution $z:=1-x^{2}$, and the fact that $\frac{\alpha}{r}<\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$ to estimate the integral in $z$.
We obtain
$I_{1}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\,d\nu\,dv=\|f_{l}\|_{L^{q}_{v}}^{q}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
Finally, to estimate $I_{3}$, we use the fact that $|v|\leq E^{1/2}$, followed by (2.12) and (2.28) to bound
$I_{3}$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{\frac{1-\alpha}{2}}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{E>16}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}>E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$\approx\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{\frac{1-\alpha}{2}}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{\left(1-\widehat{u}\cdot\sigma\right)^{\alpha}}\mathds{1}_{E>16}\mathds{1}_{-1\leq\widehat{u}\cdot\sigma<1-2E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\mathds{1}_{E>16}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{-1\leq\widehat{u}\cdot\sigma\leq 1-2E^{-1/2}}}{\left(1-\widehat{u}\cdot\sigma\right)^{\alpha}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\mathds{1}_{E>16}\int_{-1}^{1-2E^{-1/2}}\frac{1}{\left(1-x\right)^{\alpha}}\,dx\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\mathds{1}_{E>16}\int_{2E^{-1/2}}^{2}z^{-\alpha}\,dz\,dv_{1}\,dv$
$\lesssim\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{q}}^{q}\|g_{k}\|_{L_{v}^{p}}^{p},$
where we used the fact that $\alpha>1$ to estimate the integral in $z$.
Combining (4.16) with the bounds on $I_{1},I_{2},I_{3}$, we deduce
$|I(\phi)|\lesssim\|f_{l}\|_{L^{q}_{v}}\|g_{k}\|_{L^{p}_{v}}\|\phi\|_{L^{r^{\prime}}_{v}}.$
Since $\phi\in L^{r^{\prime}}_{v}$ was chosen arbitrarily, we conclude
(4.17) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{s,1}^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
Estimate for $Q_{s,2}^{+}$
Again, we argue by duality. Fixing $\phi\in L_{v}^{r^{\prime}}$, we estimate the quantity
$J(\phi):=\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{k+1-\gamma}Q_{s,2}^{+}(v)\phi(v)\,dv.$
Using (4.7), the exponents’ compatibility condition, and Hölder’s inequality, we obtain
$|J(\phi)|$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}^{\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}|f_{l}(v^{*})|^{p/q^{\prime}}|\phi(v)|^{r^{\prime}/q^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p^{\prime}}}\right)$
$\hskip 85.35826pt\times\left(\frac{{\langle}v{\rangle}^{2/p^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p^{\prime}}|g_{k}(v_{1}^{*})|^{q/p^{\prime}}|\phi(v)|^{r^{\prime}/p^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left({\langle}v{\rangle}^{1/r}|f_{l}(v^{*})|^{p/r}|g_{k}(v_{1}^{*})|^{q/r}|\right)\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}\,dv$
(4.18) $\leq J_{1}^{1/q^{\prime}}J_{2}^{1/p^{\prime}}J_{3}^{1/r},$
where
$J_{1}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-\frac{q^{\prime}}{p^{\prime}}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{q^{\prime}(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q^{\prime}/p^{\prime}}}|f_{l}(v^{*})|^{p}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}\,dv,$
$J_{2}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp^{\prime}}}|g_{k}(v_{1}^{*})|^{q}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}\,dv,$
$J_{3}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v{\rangle}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}\,dv.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $p^{\prime}l>q^{\prime}l>2$, we use (2.26) to bound
$J_{2}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\left(\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp^{\prime}}}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{|v_{1}^{*}|^{2}>2}\,d\sigma\,dv_{1}\right)\,dv\lesssim\|g_{k}\|_{L^{q}_{v}}^{q}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
For $J_{1}$, since $q^{\prime}(k-1)\geq 2$, we argue similarly as in the estimate of $J_{1}$ in the proof of Proposition 4.1, to obtain
$J_{1}$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{1-\frac{q^{\prime}}{p^{\prime}}}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q^{\prime}/p^{\prime}}}\,dx\,d\nu\,dv.$
The inner integral now can be estimated as follows
$\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p/q}}\,dx\lesssim\int_{0}^{1}\frac{\mathds{1}_{1-x<{\langle}v{\rangle}^{-1}}}{(1-x)^{q^{\prime}/p^{\prime}}}\,dx\approx\int_{0}^{{\langle}v{\rangle}^{-1}}z^{-q^{\prime}/p^{\prime}}\,dz\lesssim{\langle}v{\rangle}^{\frac{q^{\prime}}{p^{\prime}}-1},$
where we used the fact that $p<q$ for the convergence of the integral in $z$. We obtain
$J_{1}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\,d\nu\,dv=\|f_{l}\|_{L^{p}_{v}}^{p}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
To estimate $J_{3}$, we note that in the corresponding domain of integration we have $|v|\gtrsim E^{1/2}$. Indeed, by (2.12) we have $|\widehat{R}_{\sigma}^{-}\cdot\sigma|<{\langle}v{\rangle}^{-1/2}<1/2$, since $|v|>4$. Now, by (2.11), (1.6), we have
$|R_{\sigma}^{-}(u)|=|u||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<\frac{|u|}{2}=\frac{|v^{*}-v_{1}^{*}|}{2}\leq\frac{|v_{1}^{*}|}{2}+\frac{|v^{*}|}{2}\leq\frac{|v_{1}^{*}|}{2}+\frac{|v|}{4},$
since $|v^{*}|<|v|/2$. Then, using (2.7) and the triangle inequality, we obtain
$|v|\geq|v_{1}^{*}|-|R_{\sigma}^{-}(u)|\geq\frac{|v_{1}^{*}|}{2}-\frac{|v|}{4}\Rightarrow|v|\geq\frac{2|v_{1}^{*}|}{5}>\frac{2E^{1/2}}{5\sqrt{2}}.$
Therefore, using the above estimate, the fact that $|v|\leq E^{1/2}$, and (2.12), (2.28), we can bound
$J_{3}$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{E>16}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}<E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{E>16}\mathds{1}_{\widehat{u}\cdot\sigma>1-2E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p}|g_{k}(v_{1})|^{q}\mathds{1}_{E>16}\int_{{\mathbb{S}}^{2}}\mathds{1}_{\widehat{u}\cdot\sigma>1-2E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$\approx\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p}|g_{k}(v_{1})|^{q}\mathds{1}_{E>16}\int_{1-2E^{-1/2}}^{1}\,dx\,dv_{1}\,dv$
$\approx\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{p}|g_{k}(v_{1})|^{q}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{p}}^{p}\|g_{k}\|_{L_{v}^{q}}^{q}.$
Combining (4.18) with the bounds on $J_{1},J_{2},J_{3}$, we deduce
$|J(\phi)|\lesssim\|f_{l}\|_{L^{p}_{v}}\|g_{k}\|_{L^{q}_{v}}\|\phi\|_{L^{r^{\prime}}_{v}}.$
Since $\phi\in L^{r^{\prime}}_{v}$ was chosen arbitrarily, we conclude
(4.19) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{s,2}^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}.$
Conclusion of the proof
Combining (4.2) with (4.13), (4.17), (4.19), estimate (4.12) follows.
4.3. Proof of Theorem 1.1
We conclude with the proof of Theorem 1.1. Using Propositions 4.14.2, we obtain
(4.20) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in C_{c}({\mathbb{R}}^{3}).$
Recalling the definition of $\|\cdot\|_{\bm{\mathcal{X}_{p,q}^{k}}}$ from (1.17), the above bound implies
(4.21) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}\|g\|_{\bm{\mathcal{X}_{p,q}^{k}}},\quad\forall f,g\in C_{c}({\mathbb{R}}^{3}).$
By Lemma A.1, we can uniquely extend $Q^{+}$ by density to a bounded bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k-1+\gamma}L^{r}_{v}.$
Finally, we prove that (1.19) holds for any $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Indeed, consider $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Using Lemma A.1, we choose sequences $\{f_{n}\}_{n=1}^{\infty}$, $\{g_{n}\}_{n=1}^{\infty}\subset C_{c}^{\infty}({\mathbb{R}}^{3})$ such that $f_{n}\to f$ and $g_{n}\to g$ as $n\to\infty$ in $\bm{\mathcal{X}_{p,q}^{k}}$. Then, by (4.20), for any $n\in\mathbb{N}$ we have
$\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f_{n},g_{n})\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{p}}.$
Letting $n\to\infty$ in the above bound, and using continuity of $Q^{+}$ for the left-hand side as well as (1.18) for the right-hand side, estimate (1.19) follows. The proof is complete.
Appendix A
In this appendix, we record a basic density result.
Lemma A.1.
Let $k\geq 0$ and $1\leq p,q<\infty$. Then $C_{c}^{\infty}({\mathbb{R}}^{3})$ is a dense subspace of $\bm{\mathcal{X}_{p,q}^{k}}$.
Proof.
Let $f\in\bm{\mathcal{X}_{p,q}^{k}}$ and $\epsilon>0$. We will construct $\psi\in C_{c}^{\infty}({\mathbb{R}}^{3})$ with $\|\psi-f\|_{\bm{\mathcal{X}_{p,q}^{k}}}<\epsilon$.
Define $\widetilde{f}={\langle}v{\rangle}^{k}f\in L^{p}_{v}\cap L^{q}_{v}$, and given $R>0$ let us write $\widetilde{f}_{R}:=\widetilde{f}\,\mathds{1}_{|x|\leq R}$. Since $\widetilde{f}\in L^{p}_{v}\cap L^{q}_{v}$, the dominated convergence theorem implies that $\widetilde{f}_{R}\to\widetilde{f}$ as $R\to\infty$ both in $L^{p}_{v}$ and $L^{q}_{v}$, thus there exists $R_{0}>0$ such that
(A.1) $\|\widetilde{f}_{R_{0}}-\widetilde{f}\|_{L^{p}_{v}}<\frac{\epsilon}{2},\quad\|\widetilde{f}_{R_{0}}-\widetilde{f}\|_{L^{q}_{v}}<\frac{\epsilon}{2}.$
Now consider a mollifier $(\rho_{\delta})_{0<\delta\leq 1}$. Then, $\rho_{\delta}\ast\widetilde{f}_{R_{0}}\in C_{c}^{\infty}({\mathbb{R}}^{3})$ and $\rho_{\delta}\ast\widetilde{f}_{R_{0}}\to\widetilde{f}_{R_{0}}$ as $\delta\to 0$ both in $L^{p}_{v}$ and $L^{q}_{v}$. Thus there exists $0<\delta_{0}<1$ such that
(A.2) $\|\rho_{\delta_{0}}\ast\widetilde{f}_{R_{0}}-\widetilde{f}_{R_{0}}\|_{L^{p}_{v}}<\frac{\epsilon}{2},\quad\|\rho_{\delta_{0}}\ast\widetilde{f}_{R_{0}}-\widetilde{f}_{R_{0}}\|_{L^{q}_{v}}<\frac{\epsilon}{2}.$
Define $\widetilde{\psi}:=(\rho_{\delta_{0}}\ast\widetilde{f}_{R_{0}})\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Then by the triangle inequality and (A.1), (A.2), we obtain
$\|\widetilde{\psi}-\widetilde{f}\|_{L^{p}_{v}}<\epsilon,\quad\|\widetilde{\psi}-\widetilde{f}\|_{L^{q}_{v}}<\epsilon.$
Defining $\psi:={\langle}v{\rangle}^{-k}\widetilde{\psi}\in C_{c}^{\infty}({\mathbb{R}}^{3})$, we have
$\|\psi-f\|_{\bm{\mathcal{X}_{p,q}^{k}}}$ $=\max\left\{\|{\langle}v{\rangle}^{k}(\psi-f)\|_{L^{p}_{v}},\|{\langle}v{\rangle}^{k}(\psi-f)\|_{L^{q}_{v}}\right\}$
$=\max\left\{\|\widetilde{\psi}-\widetilde{f}\|_{L^{p}_{v}},\|\widetilde{\psi}-\widetilde{f}\|_{L^{q}_{v}}\right\}<\epsilon,$
and the claim follows. ∎